THREE DIALOGUES

BETWEEN HYLAS AND PHILONOUS

IN OPPOSITION TO SCEPTICS AND ATHEISTS George Berkeley

Edited by David R. Wilkins

NOTE ON THE TEXT

This edition is based on the edition of the Three Dialogue between Hylas and Philonous published with the Treatise concerning the principles of human knowledge by Jacob Tonson in 1734, and follows that edition in spelling, capitalization and punctuation. However a small number of changes in punctuation have been introduced: most of these involve the replacement of a full stop by a question mark after a number of direct questions.

David R. Wilkins Dublin, December 2002

THREE

DIALOGUES

BETWEEN

Hylas and Philonous.

In OPPOSITION to

SCEPTICS and ATHEISTS.

GEORGE BERKELEY, M. A.

Fellow of Trinity College, Dublin.

First Printed in the Year 1713.

THE FIRST

DIALOGUE.

Philonous. Good Morrow, Hylas: I did not expect to find you abroad so early.

Hylas. It is indeed something unusual; but my Thoughts were so taken up with a Subject I was discoursing of last Night, that finding I could not sleep, I resolved to rise and take a turn in the Garden.

Phil. It happened well, to let you see what innocent and agreeable Pleasures you lose every Morning. Can there be a pleasanter time of the Day, or a more delightful Season of the Year? That purple Sky, these wild but sweet Notes of Birds, the fragrant Bloom upon the Trees and Flowers, the gentle Influence of the rising Sun, these and a thousand nameless Beauties of Nature inspire the Soul with secret Transports; its Faculties too being at this time fresh and lively, are fit for those Meditations, which the Solitude of a Garden and Tranquillity of the Morning naturally dispose us to. But I am afraid I interrupt your Thoughts: for you seemed very intent on something.

Hyl. It is true, I was, and shall be obliged to you if you will permit me to go on in the same Vein; not that I would by any means deprive my self of your Company, for my Thoughts always flow more easily in Conversation with a Friend, than when I am alone: But my Request is, that you would suffer me to impart my Reflexions to you.

Phil. With all my heart, it is what I should have requested my self, if you had not prevented me.

Hyl. I was considering the odd Fate of those Men who have in all Ages, through an Affectation of being distinguished from the Vulgar, or some unaccountable Turn of Thought, pretended either to believe nothing at all, or to believe the most extravagant Things in the World. This however might be born, if their Paradoxes and Scepticism did not draw after them some Consequences of general Disadvantage to Mankind. But the Mischief lieth here; that when Men of less Leisure see them who are supposed to have spent their whole time in the Pursuits of Knowledge, professing an intire Ignorance of all Things, or advancing such Notions as are repugnant to plain and commonly received Principles, they will be tempted to entertain Suspicions concerning the most important Truths, which they had hitherto held sacred and unquestionable.

Phil. I intirely agree with you, as to the ill Tendency of the affected Doubts of some Philosophers, and fantastical Conceits of others. I am even so far gone of late in this way of Thinking, that I have quitted several of the sublime Notions I had got in their Schools for vulgar Opinions. And I give it you on my Word, since this Revolt from Metaphysical Notions to the plain Dictates of Nature and common Sense, I find my Understanding strangely enlightened, so that I can now easily comprehend a great many Things which before were all Mystery and Riddle.

Hyl. I am glad to find there was nothing in the Accounts I heard of you.

Phil. Pray, what were those?

Hyl. You were represented in last Night’s Conversation, as one who maintained the most extravagant Opinion that ever entered into the Mind of Man, to wit, That there is no such Thing as material Substance in the World.

Phil. That there is no such Thing as what Philosophers call Material Substance, I am seriously persuaded: But if I were made to see any thing absurd or sceptical in this, I should then have the same Reason to renounce this, that I imagine I have now to reject the contrary Opinion.

Hyl. What! can any Thing be more fantastical, more repugnant to common Sense, or a more manifest Piece of Scepticism, than to believe there is no such Thing as Matter?

Phil. Softly, good Hylas. What if it should prove, that you, who hold there is, are by virtue of that Opinion a greater Sceptic, and maintain more Paradoxes and Repugnances to common Sense, than I who believe no such Thing?

Hyl. You may as soon persuade me, The Part is greater than the Whole, as that, in order to avoid Absurdity and Scepticism, I should ever be obliged to give up my Opinion in this Point.

Phil. Well then, are you content to admit that Opinion for true, which upon Examination shall appear most agreeable to common Sense, and remote from Scepticism?

Hyl. With all my Heart. Since you are for raising Disputes about the plainest Things in Nature, I am content for once to hear what you have to say.

Phil. Pray, Hylas, what do you mean by a Sceptic?

Hyl. I mean what all Men mean, one that doubts of every Thing.

Phil. He then who entertains no Doubt concerning some particular Point, with regard to that Point cannot be thought a Sceptic.

Hyl. I agree with you.

Phil. Whether doth Doubting consist in embracing the Affirmative or Negative Side of a Question?

Hyl. In neither; for whoever understands English, cannot but know that Doubting signifies a Suspense between both.

Phil. He then that denieth any Point, can no more be said to doubt of it, than he who affirmeth it with the same Degree of Assurance.

Hyl. True.

Phil. And consequently, for such his Denial is no more to be esteemed a Sceptic than the other.

Hyl. I acknowledge it.

Phil. How cometh it to pass then, Hylas, that you pronounce me a Sceptic, because I deny what you affirm, to wit, the Existence of Matter? Since, for aught you can tell, I am as peremptory in my Denial, as you in your Affirmation.

Hyl. Hold, Philonous, I have been a little out in my Definition; but every false Step a Man makes in Discourse is not to be insisted on. I said indeed, that a Sceptic was one who doubted of every Thing; but I should have added, or who denies the Reality and Truth of Things.

Phil. What Things? Do you mean the Principles and Theoremes of Sciences? But these you know are universal intellectual Notions, and consequently independent of Matter; the Denial therefore of this doth not imply the denying them.

Hyl. I grant it. But are there no other Things? What think you of distrusting the Senses, of denying the real Existence of sensible Things, or pretending to know nothing of them. Is not this sufficient to denominate a Man a Sceptic?

Phil. Shall we therefore examine which of us it is that denies the Reality of Sensible Things, or professes the greatest Ignorance of them; since, if I take you rightly, he is to be esteemed the greatest Sceptic?

Hyl. That is what I desire.

Phil. What mean you by Sensible Things?

Hyl. Those Things which are perceived by the Senses. Can you imagine that I mean any thing else?

Phil. Pardon me, Hylas, if I am desirous clearly to apprehend your Notions, since this may much shorten our Inquiry. Suffer me then to ask you this farther Question. Are those Things only perceived by the Senses which are perceived immediately? Or may those Things properly be said to be Sensible, which are perceived mediately, or not without the Intervention of others?

Hyl. I do not sufficiently understand you.

Phil. In reading a Book, what I immediately perceive are the Letters, but mediately, or by means of these, are suggested to my Mind the notions of God, Virtue, Truth, &c. Now, that the Letters are truly Sensible Things, or perceived by Sense, there is no doubt: But I would know whether you take the Things suggested by them to be so too.

Hyl. No certainly, it were absurd to think God or Virtue Sensible Things, though they may be signified and suggested to the Mind by Sensible Marks, with which they have an arbitrary Connexion.

Phil. It seems then, that by Sensible Things you mean those only which can be perceived immediately by Sense.

Hyl. Right.

Phil. Doth it not follow from this, that though I see one part of the Sky Red, and another Blue, and that my Reason doth thence evidently conclude there must be some Cause of that Diversity of Colours, yet that Cause cannot be said to be a Sensible Thing, or perceived by the Sense of Seeing?

Hyl. It doth.

Phil. In like manner, though I hear Variety of Sounds, yet I cannot be said to hear the Causes of those Sounds.

Hyl. You cannot.

Phil. And when by my Touch I perceive a Thing to be hot and heavy, I cannot say with any Truth or Propriety, that I feel the Cause of its Heat or Weight.

Hyl. To prevent any more Questions of this kind, I tell you once for all, that by Sensible Things I mean those only which are perceived by Sense, and that in truth the Senses perceive nothing which they do not perceive immediately: for they make no Inferences. The deducing therefore of Causes or Occasions from Effects and Appearances, which alone are perceived by Sense, intirely relates to Reason.

Phil. This Point then is agreed between us, That Sensible Things are those only which are immediately perceived by Sense. You will farther inform me, whether we immediately perceive by Sight any thing beside Light, and Colours, and Figures: or by Hearing, any thing but Sounds: by the Palate, any thing beside Tastes: by the Smell, beside Odors: or by the Touch, more than tangible Qualities.

Hyl. We do not. Phil. It seems therefore, that if you take away all sensible Qualities, there remains

nothing sensible.

Hyl. I grant it. Phil. Sensible Things therefore are nothing else but so many sensible Qualities, or

Combinations of sensible Qualities.

Hyl. Nothing else. Phil. Heat then is a sensible Thing. Hyl. Certainly. Phil. Doth the Reality of sensible Things consist in being perceived? or, is it something

distinct from their being perceived, and that bears no relation to the Mind?

Hyl. To exist is one thing, and to be perceived is another. Phil. I speak with regard to sensible Things only: And of these I ask, Whether by their

real Existence you mean a Subsistence exterior to the Mind, and distinct from their being perceived?

Hyl. I mean a real absolute Being, distinct from, and without any relation to their being perceived.

Phil. Heat therefore, if it be allowed a real Being, must exist without the Mind. Hyl. It must. Phil. Tell me, Hylas, is this real Existence equally compatible to all Degrees of Heat,

which we perceive: or is there any Reason why we should attribute it to some, and deny it others? And if there be, pray let me know that Reason.

Hyl. Whatever Degree of Heat we perceive by Sense, we may be sure the same exists in the Object that occasions it.

Phil. What, the greatest as well as the least? Hyl. I tell you, the Reason is plainly the same in respect of both: They are both perceived

by Sense; nay, the greater Degree of Heat is more sensibly perceived; and consequently, if there is any Difference, we are more certain of its real Existence than we can be of the Reality of a lesser Degree.

Phil. But is not the most vehement and intense Degree of Heat a very great Pain? Hyl. No one can deny it. Phil. And is any unperceiving Thing capable of Pain or Pleasure? Hyl. No certainly. Phil. Is your material Substance a senseless Being, or a Being endowed with Sense and

Perception?

Hyl. It is senseless, without doubt. Phil. It cannot therefore be the Subject of Pain. Hyl. By no means. Phil. Nor consequently of the greatest Heat perceived by Sense, since you acknowledge

this to be no small Pain.

Hyl. I grant it. Phil. What shall we say then of your external Object; is it a material Substance, or no? Hyl. It is a material Substance with the sensible Qualities inhering in it.

Phil. How then can a great Heat exist in it, since you own it cannot in a material Substance? I desire you would clear this Point.

Hyl. Hold, Philonous. I fear I was out in yielding intense Heat to be a Pain. It should seem rather, that Pain is something distinct from Heat, and the Consequence or Effect of it.

Phil. Upon putting your Hand near the Fire, do you perceive one simple uniform Sensation, or two distinct Sensations?

Hyl. But one simple Sensation. Phil. Is not the Heat immediately perceived? Hyl. It is. Phil. And the Pain? Hyl. True. Phil. Seeing therefore they are both immediately perceived at the same time, and the

Fire affects you only with one simple, or uncompounded Idea, it follows that this same simple Idea is both the intense Heat immediately perceived, and the Pain; and consequently, that the intense Heat immediately perceived, is nothing distinct from a particular sort of Pain.

Hyl. It seems so. Phil. Again, try in your Thoughts, Hylas, if you can conceive a vehement Sensation to

be without Pain, or Pleasure.

Hyl. I cannot. Phil. Or can you frame to yourself an Idea of sensible Pain or Pleasure in general,

abstracted from every particular Idea of Heat, Cold, Tastes, Smells? &c.

Hyl. I do not find that I can. Phil. Doth it not therefore follow, that sensible Pain is nothing distinct from those

Sensations or Ideas, in an intense Degree?

Hyl. It is undeniable; and to speak the Truth, I begin to suspect a very great Heat cannot exist but in a Mind perceiving it.

Phil. What! are you then in that Sceptical State of Suspense, between Affirming and Denying?

Hyl. I think I may be positive in the Point. A very violent and painful Heat cannot exist without the Mind.

Phil. It hath not therefore, according to you, any real Being. Hyl. I own it. Phil. Is it therefore certain, that there is no body in Nature really hot? Hyl. I have not denied there is any real Heat in Bodies. I only say, there is no such thing

as an intense real Heat.

Phil. But did you not say before, that all Degrees of Heat were equally real: or if there was any difference, that the Greater were more undoubtedly real than the Lesser?

Hyl. True: But it was, because I did not then consider the Ground there is for distinguishing between them, which I now plainly see. And it is this: Because intense Heat is nothing else but a particular kind of painful Sensation; and Pain cannot exist but in a perceiving Being; it follows that no intense Heat can really exist in an unperceiving corporeal Substance. But this is no Reason why we should deny Heat in an inferior Degree to exist in such a Substance.

Phil. But how shall we be able to discern those Degrees of Heat which exist only in the Mind, from those which exist without it?

Hyl. That is no difficult matter. You know, the least Pain cannot exist unperceived; whatever therefore Degree of Heat is a Pain, exists only in the Mind. But as for all other Degrees of Heat, nothing obliges us to think the same of them.

Phil. I think you granted before, that no unperceiving Being was capable of Pleasure, any more than of Pain.

Hyl. I did. Phil. And is not Warmth, or a more gentle Degree of Heat than what causes Uneasiness,

a Pleasure?

Hyl. What then? Phil. Consequently it cannot exist without the Mind in any unperceiving Substance, or

Body.

Hyl. So it seems. Phil. Since therefore, as well as those Degrees of Heat that are not painful, as those that

are, can exist only in a Thinking Substance; may we not conclude that external Bodies are absolutely incapable of any Degree of Heat whatsoever?

Hyl. On second Thoughts, I do not think it so evident that Warmth is a Pleasure, as that a great Degree of Heat is a Pain.

Phil. I do not pretend that Warmth is as great a Pleasure as Heat is a Pain. But if you grant it to be even a small Pleasure, it serves to make good my Conclusion.

Hyl. I could rather call it an Indolence. It seems to be nothing more than a Privation of both Pain and Pleasure. And that such a Quality or State as this may agree to an unthinking Substance, I hope you will not deny.

Phil. If you are resolved to maintain that Warmth, or a gentle Degree of Heat, is no Pleasure, I know not how to convince you otherwise, than by appealing to your own Sense. But what think you of Cold?

Hyl. The same that I do of Heat. An intense Degree of Cold is a Pain; for to feel a very great Cold, is to perceive a great Uneasiness: It cannot therefore exist without the Mind; but a lesser Degree of Cold may, as well as a lesser Degree of Heat.

Phil. Those Bodies therefore, upon whose Application to our own, we perceive a moderate Degree of Heat, must be concluded to have a moderate Degree of Heat or Warmth in them; And those, upon whose Application we feel a like Degree of Cold, must be thought to have Cold in them.

Hyl. They must. Phil. Can any Doctrine be true that necessarily leads a Man into an Absurdity? Hyl. Without doubt it cannot. Phil. Is it not an Absurdity to think that the same thing should be at the same time

both cold and warm?

Hyl. It is. Phil. Suppose now one of your Hands hot, and the other cold, and that they are both

at once put into the same Vessel of Water, in an intermediate State; will not the Water seem cold to one Hand, and warm to the other?

Hyl. It will. Phil. Ought we not therefore by your Principles to conclude, it is really both cold and

warm at the same time, that is, according to your own Concession, to believe an Absurdity. Hyl. I confess it seems so.

Phil. Consequently, the Principles themselves are false, since you have granted that no true Principle leads to an Absurdity.

Hyl. But after all, can any thing be more absurd than to say, there is no Heat in the Fire?

Phil. To make the Point still clearer; tell me, whether in two Cases exactly alike, we ought not to make the same Judgment?

Hyl. We ought.

Phil. When a Pin pricks your Finger, doth it not rend and divide the Fibres of your Flesh?

Hyl. It doth.

Phil. And when a Coal burns your Finger, doth it any more?

Hyl. It doth not.

Phil. Since therefore you neither judge the Sensation itself occasioned by the Pin, nor any thing like it to be in the Pin; you should not, conformably to what you have now granted, judge the Sensation occasioned by the Fire, or any thing like it, to be in the Fire.

Hyl. Well, since it must be so, I am content to yield this Point, and acknowledge, that Heat and Cold are only Sensations existing in our Minds: But there still remain Qualities enough to secure the Reality of external Things.

Phil. But what will you say, Hylas, if it shall appear that the Case is the same with regard to all other sensible Qualities, and that they can no more be supposed to exist without the Mind, than Heat or Cold?

Hyl. Then indeed you will have done something to the purpose; but that is what I despair of seeing proved.

Phil. Let us examine them in order. What think you of Tastes, do they exist without the Mind, or no?

Hyl. Can any Man in his Senses doubt whether Sugar is sweet, or Wormwood bitter? Phil. Inform me, Hylas. Is a sweet Taste a particular kind of Pleasure or pleasant

Sensation, or is it not?

Hyl. It is.

Phil. And is not Bitterness some kind of Uneasiness or Pain?

Hyl. I grant it.

Phil. If therefore Sugar and Wormwood are unthinking corporeal Substances existing without the Mind, how can Sweetness and Bitterness, that is, Pleasure and Pain, agree to them?

Hyl. Hold, Philonous, I now see what it was deluded me all this time. You asked whether Heat and Cold, Sweetness and Bitterness, were not particular Sorts of Pleasure and Pain; to which I answered simply, that they were. Whereas I should have thus distinguished: Those Qualities, as perceived by us, are Pleasures or Pains, but not as existing in the external Objects. We must not therefore conclude absolutely, that there is no Heat in the Fire, or Sweetness in the Sugar, but only that Heat or Sweetness, as perceived by us, are not in the Fire or Sugar. What say you to this?

Phil. I say it is nothing to the Purpose. Our Discourse proceeded altogether concerning Sensible Things, which you defined to be the Things we immediately perceive by our Senses. Whatever other Qualities therefore you speak of, as distinct from these, I know nothing of them, neither do they at all belong to the Point in Dispute. You may indeed pretend to have discovered certain Qualities which you do not perceive, and assert those insensible Qualities exist in Fire and Sugar. But what Use can be made of this to your present Purpose, I am at a loss to conceive. Tell me then once more, do you acknowledge that Heat and Cold, Sweetness and Bitterness, (meaning those Qualities which are perceived by the Senses) do not exist without the Mind?

Hyl. I see it is to no purpose to hold out, so I give up the Cause as to those mentioned Qualities. Though I profess it sounds odly, to say that Sugar is not sweet.

Phil. But for your farther Satisfaction, take this along with you: That which at other times seems sweet, shall to a distempered Palate appear bitter. And nothing can be plainer, than that divers Persons perceive different Tastes in the same Food, since that which one Man delights in, another abhors. And how could this be, if the Taste was something really inherent in the Food?

Hyl. I acknowledge I know not how.

Phil. In the next place, Odours are to be considered. And with regard to these, I would fain know, whether what hath been said of Tastes doth not exactly agree to them? Are they not so many pleasing or displeasing Sensations?

Hyl. They are.

Phil. Can you then conceive it possible that they should exist in an unperceiving Thing? Hyl. I cannot. Phil. Or can you imagine, that Filth and Ordure affect those brute Animals that feed

on them out of Choice, with the same Smells which we perceive in them?

Hyl. By no means.

Phil. May we not therefore conclude of Smells, as of the other forementioned Qualities, that they cannot exist in any but a perceiving Substance or Mind?

Hyl. I think so.

Phil. Then as to Sounds, what must we think of them: Are they Accidents really inherent in external Bodies, nor not?

Hyl. That they inhere not in the sonorous Bodies, is plain from hence; because a Bell struck in the exhausted Receiver of an Air-Pump, sends forth no Sound. The Air therefore must be thought the Subject of Sound.

Phil. What Reason is there for that, Hylas?

Hyl. Because when any Motion is raised in the Air, we perceive a Sound greater or lesser, in Proportion to the Air’s Motion; but without some Motion in the Air, we never hear any Sound at all.

Phil. And granting that we never hear a Sound but when some Motion is produced in the Air, yet I do not see how you can infer from thence, that the Sound itself is in the Air.

Hyl. It is this very Motion in the external Air, that produces in the Mind the Sensation of Sound. For, striking on the Drum of the Ear, it causeth a Vibration, which by the Auditory Nerves being communicated to the Brain, the Soul is thereupon affected with the Sensation called Sound.

Phil. What! is Sound a Sensation?

Hyl. I tell you, as perceived by us, it is a particular Sensation in the Mind.

Phil. And can any Sensation exist without the Mind?

Hyl. No certainly.

Phil. How then can Sound, being a Sensation exist in the Air, if by the Air you mean a senseless Substance existing without the Mind?

Hyl. You must distinguish, Philonous, between Sound as it is perceived by us, and as it is in itself; or (which is the same thing) between the Sound we immediately perceive, and that which exists without us. The former indeed is a particular kind of Sensation, but the latter is merely a Vibrative or Undulatory Motion in the Air.

Phil. I thought I had already obviated that Distinction by the Answer I gave when you were applying it in a like Case before. But to say no more of that; Are you sure then that Sound is really nothing but Motion?

Hyl. I am.

Phil. Whatever therefore agrees to real Sound, may with Truth be attributed to Motion. Hyl. It may. Phil. It is then good Sense to speak of Motion, as of a thing that is loud, sweet, acute,

or grave.

Hyl. I see you are resolved not to understand me. Is it not evident, those Accidents or Modes belong only to sensible Sound, or Sound in the common Acceptation of the Word, but not to Sound in the Real and Philosophic Sense, which, as I just now told you, is nothing but a certain Motion of the Air?

Phil. It seems then there are two Sorts of Sound, the one Vulgar, or that which is heard, the other Philosophical and Real.

Hyl. Even so.

Phil. And the latter consists in Motion.

Hyl. I told you so before.

Phil. Tell me, Hylas, to which of the Senses think you, the Idea of Motion belongs: To the Hearing?

Hyl. No certainly, but to the Sight and Touch.

Phil. It should follow then, that according to you, real Sounds may possibly be seen or felt, but never heard.

Hyl. Look you, Philonous, you may, if you please make a Jest of my Opinion but that will not alter the Truth of Things. I own indeed, the Inferences you draw me into, sound something odly; but common Language, you know, is framed by, and for the Use of the Vulgar: we must not therefore wonder, if Expressions adapted to exact Philosophic Notions, seem uncouth and out of the way.

Phil. Is it come to that? I assure you, I imagine myself to have gained no small Point, since you make so light of departing from common Phrases and Opinions; it being a main Part of our Inquiry, to examine whose Notions are widest of the common Road, and most repugnant to the general Sense of the World. But can you think it no more than a Philosophical Paradox, to say that real Sounds are never heard, and that the Idea of them is obtained by some other Sense. And is there nothing in this contrary to Nature and the Truth of Things?

Hyl. To deal ingenuously, I do not like it. And after the Concessions already made, I had as well grant that Sounds too have no real Being without the Mind.

Phil. And I hope you will make no Difficulty to acknowledge the same of Colours.

Hyl. Pardon me: the Case of Colours is very different. Can any thing be plainer, than that we see them on the Objects?

Phil. The Objects you speak of are, I suppose, corporeal Substances existing without the Mind.

Hyl. They are. Phil. And have true and real Colours inhering in them? Hyl. Each visible Object hath that Colour which we see in it. Phil. How! Is there any thing visible but what we perceive by Sight? Hyl. There is not. Phil. And do we perceive any thing by Sense, which we do not perceive immediately? Hyl. How often must I be obliged to repeat the same thing? I tell you, we do not. Phil. Have Patience, good Hylas; and tell me once more, whether there is any thing

immediately perceived by the Senses, except sensible Qualities. I know you asserted there was not: But I would now be informed, whether you still persist in the same Opinion.

Hyl. I do. Phil. Pray, is your corporeal Substance either a sensible Quality, or made up of sensible

Qualities?

Hyl. What a Question that is! who ever thought it was? Phil. My Reason for asking was, because in saying, each visible Object hath that Colour

which we see in it, you make visible Objects to be corporeal Substances; which implies either that corporeal Substances are sensible Qualities, or else that there is something beside sensible Qualities perceived by Sight: But as this Point was formerly agreed between us, and is still maintained by you, it is a clear Consequence, that your corporeal Substance is nothing distinct from sensible Qualities.

Hyl. You may draw as many absurd Consequences as you please, and endeavour to perplex the plainest Things; but you shall never persuade me out of my Senses. I clearly understand my own Meaning.

Phil. I wish you would make me understand it too. But since you are unwilling to have your Notion of corporeal Substance examined, I shall urge that Point no farther. Only be pleased to let me know, whether the same Colours which we see, exist in external Bodies, or some other.

Hyl. The very same. Phil. What! are then the beautiful Red and Purple we see on yonder Clouds, really in

them? Or do you imagine they have in themselves any other Form, than that of a dark Mist or Vapour?

Hyl. I must own, Philonous, those Colours are not really in the Clouds as they seem to be at this Distance. They are only apparent Colours.

Phil. Apparent call you them? how shall we distinguish these apparent Colours from real?

Hyl. Very easily. Those are to be thought apparent, which appearing only at a distance, vanish upon a nearer Approach.

Phil. And those I suppose are to be thought real, which are discovered by the most near and exact Survey.

Hyl. Right. Phil. Is the nearest and exactest Survey made by the help of a Microscope, or by the

naked Eye?

Hyl. By a Microscope, doubtless.

Phil. But a Microscope often discovers Colours in an Object different from those perceived by the unassisted Sight. And in case we had Microscopes magnifying to any assigned Degree; it is certain, that no Object whatsoever viewed through them, would appear in the same Colour which it exhibits to the naked Eye.

Hyl. And what will you conclude from all this? You cannot argue that there are really and naturally no Colours on Objects: because by artificial Managements they may be altered, or made to vanish.

Phil. I think it may evidently be concluded from your own Concessions, that all the Colours we see with our naked Eyes, are only apparent as those on the Clouds, since they vanish upon a more close and accurate Inspection, which is afforded us by a Microscope. Then as to what you say by way of Prevention: I ask you, whether the real and natural State of an Object is better discovered by a very sharp and piercing Sight, or by one which is less sharp?

Hyl. By the former without doubt.

Phil. Is it not plain from Dioptrics, that Microscopes make the Sight more penetrating, and represents Objects as they would appear to the Eye, in case it were naturally endowed with a most exquisite Sharpness?

Hyl. It is.

Phil. Consequently the Microscopical Representation is to be thought that which best sets forth the real Nature of the Thing, or what it is in itself. The Colours therefore by it perceived, are more genuine and real, than those perceived otherwise.

Hyl. I confess there is something in what you say.

Phil. Besides, it is not only possible but manifest, that there actually are Animals, whose Eyes are by Nature framed to perceive those Things, which by reason of their Minuteness escape our Sight. What think you of those inconceivably small Animals perceived by Glasses? Must we suppose they are all stark blind? Or, in case they see, can it be imagined their Sight hath not the same Use in preserving their Bodies from Injuries, which appears in That of all other Animals? And if it hath, is it not evident, they must see Particles less than their own Bodies, which will present them with a far different View in each Object, from that which strikes our Senses? Even our own Eyes do not always represent Objects to us after the same manner. In the Jaundice, every one knows that all Things seem yellow. Is it not therefore highly probable, those Animals in whose Eyes we discern a very different Texture from that of ours, and whose Bodies abound with different Humours, do not see the same Colours in every Object that we do? From all which, should it not seem to follow, that all Colours are equally apparent, and that none of those which we perceive are really inherent in any outward Object?

Hyl. It should.

Phil. The Point will be past all doubt, if you consider, that in case Colours were real Properties or Affections inherent in external Bodies, they could admit of no Alteration, without some Change wrought in the very Bodies themselves: But is it not evident from what hath been said, that upon the Use of Microscopes, upon a Change happening in the Humours of the Eye, or a Variation of Distance, without any manner of real Alteration in the Thing itself, the Colours of any Object are either changed, or totally disappear? Nay all other Circumstances remaining the same, change but the Situation of some Objects, and they shall present different Colours to the Eye. The same thing happens upon viewing an Object in various Degrees of Light. And what is more known, than that the same Bodies appear differently coloured by Candle-light, from what they do in the open Day? Add to these the Experiment of a Prism, which separating the heterogeneous Rays of Light, alters the Colour of any Object; and will cause the Whitest to appear of a deep Blue or Red to the naked Eye. And now tell me, whether you are still of Opinion, that every Body hath its true real Colour inhering in it; and if you think it hath, I would fain know farther from you, what certain Distance and Position of the Object, what peculiar Texture and Formation of the Eye, what Degree or Kind of Light is necessary for ascertaining that true Colour, and distinguishing it from apparent ones.

Hyl. I own myself intirely satisfied, that they are all equally apparent; and that there is no such thing as Colour really inhering in external Bodies, but that it is altogether in the Light. And what confirms me in this Opinion is, that in proportion to the Light, Colours are still more or less vivid; and if there be no Light, then are there no Colours perceived. Besides, allowing there are Colours on external Objects, yet how is it possible for us to perceive them? For no external Body affects the Mind, unless it act first on our Organs of Sense. But the only Action of Bodies is Motion; and Motion cannot be communicated otherwise than by Impulse. A distant Object therefore cannot act on the Eye, nor consequently make itself or its Properties perceivable to the Soul. Whence it plainly follows, that it is immediately some contiguous Substance, which operating on the Eye occasions a Perception of Colours: And such is Light.

Phil. How! is Light then a Substance?

Hyl. I tell you, Philonous, external Light is nothing but a thin fluid Substance, whose minute Particles being agitated with a brisk Motion, and in various Manners reflected from the different Surfaces of outward Objects to the Eyes, communicate different Motions to the Optick Nerves; which being propagated to the Brain, cause therein various Impressions: And these are attended with the Sensations of Red, Blue, Yellow, &c.

Phil. It seems then, the Light doth no more than shake the Optick Nerves.

Hyl. Nothing else.

Phil. And consequent to each particular Motion of the Nerves the Mind is affected with a Sensation, which is some particular Colour.

Hyl. Right.

Phil. And these Sensations have no Existence without the Mind.

Hyl. They have not.

Phil. How then do you affirm that Colours are in the Light, since by Light you understand a corporeal Substance external to the Mind?

Hyl. Light and Colours, as immediately perceived by us, I grant cannot exist without the Mind. But in themselves they are only the Motions and Configurations of certain insensible Particles of Matter.

Phil. Colours then in the vulgar Sense, or taken for the immediate Objects of Sight, cannot agree to any but a perceiving Substance.

Hyl. That is what I say.

Phil. Well then, since you give up the point as to those sensible Qualities, which are alone thought Colours by all Mankind beside, you may hold what you please with regard to those invisible ones of the Philosophers. It is not my Business to dispute about them; only I would advise you to bethink your self, whether considering the Inquiry we are upon, it be prudent for you to affirm, the Red and Blue which we see are not real Colours, but certain unknown Motions and Figures which no Man ever did or can see, are truly so. Are not these shocking Notions, and are not they subject to as many ridiculous Inferences, as those you were obliged to renounce before in the Case of Sounds?

Hyl. I frankly own, Philonous, that it is in vain to stand out any longer. Colours, Sounds, Tastes, in a word, all those termed Secondary Qualities, have certainly no Existence without the Mind. But by this Acknowledgment I must not be supposed to derogate any thing from the Reality of Matter or external Objects, seeing it is no more than several Philosophers maintain, who nevertheless are the farthest imaginable from denying Matter. For the clearer Understanding of this, you must know sensible Qualities are by Philosophers divided into Primary and Secondary. The former are Extension, Figure, Solidity, Gravity, Motion, and Rest. And these they hold exist really in Bodies. The latter are those above enumerated; or briefly, all sensible Qualities beside the Primary, which they assert are only so many Sensations or Ideas existing no where but in the Mind. But all this, I doubt not, you are already apprised of. For my part, I have been a long time sensible there was such an Opinion current among Philosophers, but was never thoroughly convinced of its Truth till now.

Phil. You are still then of Opinion, that Extension and Figures are inherent in external unthinking Substances.

Hyl. I am.

Phil. But what if the same Arguments which are brought against Secondary Qualities, will hold proof against these also?

Hyl. Why then I shall be obliged to think, they too exist only in the Mind.

Phil. Is it your Opinion, the very Figure and Extension which you perceive by Sense, exist in the outward Object or material Substance?

Hyl. It is.

Phil. Have all other Animals as good Grounds to think the same of the Figure and Extension which they see and feel?

Hyl. Without doubt, if they have any Thought at all.

Phil. Answer me, Hylas. Think you the Senses were bestowed upon all Animals for their Preservation and Well-being in Life? or were they given to Men alone for this End?

Hyl. I make no question but they have the same Use in all other Animals.

Phil. If so, is it not necessary they should be enabled by them to perceive their own Limbs, and those Bodies which are capable of harming them?

Hyl. Certainly.

Phil. A Mite therefore must be supposed to see his own Foot, and Things equal or even less than it, as Bodies of some considerable Dimension; though at the same time they appear to you scarce discernible, or at best as so many visible Points.

Hyl. I cannot deny it.

Phil. And to Creatures less than the Mite they will seem yet larger.

Hyl. They will.

Phil. Insomuch that what you can hardly discern, will to another extremely minute Animal appear as some huge Mountain.

Hyl. All this I grant.

Phil. Can one and the same thing be at the same time in itself of different Dimensions? Hyl. That were absurd to imagine.

Phil. But from what you have laid down it follows, that both the Extension by you perceived, and that perceived by the Mite itself, as likewise all those perceived by lesser Animals, are each of them the true Extension of the Mite’s Foot, that is to say, by your own Principles you are led into an Absurdity.

Hyl. There seems to be some Difficulty in the Point.

Phil. Again, have you not acknowledged that no real inherent Property of any Object can be changed, without some Change in the thing itself?

Hyl. I have.

Phil. But as we approach to or recede from an Object, the visible Extension varies, being at one Distance ten or a hundred times greater than at another. Doth it not therefore follow from hence likewise, that it is not really inherent in the Object?

Hyl. I own I am at a loss what to think.

Phil. Your Judgment will soon be determined, if you will venture to think as freely concerning this Quality, as you have done concerning the rest. Was it not admitted as a good Argument, that neither Heat nor Cold was in the Water, because it seemed warm to one Hand, and cold to the other?

Hyl. It was.

Phil. Is it not the very same Reasoning to conclude, there is no Extension or Figure in an Object, because to one Eye it shall seem little, smooth, and round, when at the same time it appears to the other, great, uneven, and angular?

Hyl. The very same. But doth this latter Fact ever happen?

Phil. You may at any time make the Experiment, by looking with one Eye bare, and with the other through a Microscope.

Hyl. I know not how to maintain it, and yet I am loth to give up Extension, I see so many odd Consequences following upon such a Concession.

Phil. Odd, say you? After the Concessions already made, I hope you will stick at nothing for its Oddness. But on the other hand should it not seem very odd, if the general reasoning which includes all other sensible Qualities did not also include Extension? If it be allowed that no Idea nor any thing like an Idea can exist in an unperceiving Substance, then surely it follows, that no Figure or Mode of Extension, which we can either perceive or imagine, or have any Idea of, can be really inherent in Matter; not to mention the peculiar Difficulty there must be, in conceiving a material Substance, prior to and distinct from Extension, to be the Substratum of Extension. Be the sensible Quality what it will, Figure, or Sound, or Colour; it seems alike impossible it should subsist in that which doth not perceive it.

Hyl. I give up the Point for the present, reserving still a Right to retract my Opinion, in case I shall hereafter discover any false Step in my Progress to it.

Phil. That is a Right you cannot be denied. Figures and Extension being dispatched, we proceed next to Motion. Can a real Motion in any external Body be at the same time both very swift and very slow?

Hyl. It cannot.

Phil. Is not the Motion of a Body swift in a reciprocal Proportion to the time it takes up in describing any given Space? Thus a Body that describes a Mile in an Hour, moves three times faster than it would in case it described only a Mile in three Hours.

Hyl. I agree with you.

Phil. And is not Time measured by the Succession of Ideas in our Minds?

Hyl. It is.

Phil. And is it not possible Ideas should succeed one another twice as fast in your Mind, as they do in mine, or in that of some Spirit of another kind?

Hyl. I own it.

Phil. Consequently the same Body may to another seem to perform its Motion over any Space in half the time that it doth to you. And the same Reasoning will hold as to any other Proportion: That is to say, according to your Principles (since the Motions perceived are both really in the Object) it is possible one and the same Body shall be really moved the same way at once, both very swift and very slow. How is this consistent either with common Sense, or with what you just now granted?

Hyl. I have nothing to say to it.

Phil. Then as to Solidity; either you do not mean any sensible Quality by that Word, and so it is beside our Inquiry: Or if you do, it must be either Hardness or Resistance. But both the one and the other are plainly relative to our Senses: It being evident, that what seems hard to one Animal, may appear soft to another, who hath greater Force and Firmness of Limbs. Nor is it less plain, that the Resistance I feel is not in the Body.

Hyl. I own the very Sensation of Resistance, which is all you immediately perceive, is not in the Body, but the Cause of that Sensation is.

Phil. But the Causes of our Sensations are not Things immediately perceived, and therefore not sensible. This Point I thought had been already determined.

Hyl. I own it was; but you will pardon me if I seem a little embarrassed: I know not how to quit my old Notions.

Phil. To help you out, do but consider, that if Extension be once acknowledged to have no Existence without the Mind, the same must necessarily be granted of Motion, Solidity, and Gravity, since they all evidently suppose Extension. It is therefore superfluous to inquire particularly concerning each of them. In denying Extension, you have denied them all to have any real Existence.

Hyl. I wonder, Philonous, if what you say be true, why those Philosophers who deny the Secondary Qualities any real Existence, should yet attribute it to the Primary. If there is no Difference between them, how can this be accounted for?

Phil. It is not my business to account for every Opinion of the Philosophers. But among other Reasons which may be assigned for this, it seems probable, that Pleasure and Pain being rather annexed to the former than the latter, may be one. Heat and Cold, Tastes and Smells, have something more vividly pleasing or disagreeable than the Ideas of Extension, Figure, and Motion, affect us with. And it being too visibly absurd to hold, that Pain or Pleasure can be in an unperceiving Substance, Men are more easily weaned from believing the external Existence of the Secondary, than the Primary Qualities. You will be satisfied there is something in this, if you recollect the Difference you made between an intense and more moderate Degree of Heat, allowing the one a real Existence, while you denied it to the other. But after all, there is no rational Ground for that Distinction; for surely an indifferent Sensation is as truly a Sensation, as one more pleasing or painful; and consequently should not any more than they be supposed to exist in an unthinking Subject.

Hyl. It is just come into my Head, Philonous, that I have somewhere heard of a Distinction between absolute and sensible Extension. Now though it be acknowledged that great and small, consisting merely in the Relation which other extended Beings have to the Parts of our own Bodies, do not really inhere in the Substances themselves; yet nothing obliges us to hold the same with regard to absolute Extension, which is something abstracted from great and small, from this or that particular Magnitude or Figure. So likewise as to Motion, swift and slow are altogether relative to the Succession of Ideas in our own Minds. But it doth not follow, because those Modifications of Motion exist not without the Mind, that therefore absolute Motion abstracted from them doth not.

Phil. Pray what is it that distinguishes one Motion, or one Part of Extension from another? Is it not something sensible, as some Degree of Swiftness or Slowness, some certain Magnitude or Figure peculiar to each?

Hyl. I think so.

Phil. These Qualities therefore stripped of all sensible Properties, are without all specific and numerical Differences, as the Schools call them.

Hyl. They are.

Phil. That is to say, they are Extension in general, and Motion in general.

Hyl. Let it be so.

Phil. But it is a universally received Maxim, that Every thing which exists, is particular. How then can Motion in general, or Extension in general exist in any corporeal Substance?

Hyl. I will take time to solve your Difficulty.

Phil. But I think the Point may be speedily decided. Without doubt you can tell, whether you are able to frame this or that Idea. Now I am content to put our Dispute on this Issue. If you can frame in your Thoughts a distinct abstract Idea of Motion or Extension, divested of all those sensible Modes, as swift and slow, great and small, round and square, and the like, which are acknowledged to exist only in the Mind, I will then yield the Point you contend for. But if you cannot, it will be unreasonable on your Side to insist any longer upon what you have no Notion of.

Hyl. To confess ingenuously, I cannot.

Phil. Can you even separate the Ideas of Extension and Motion, from the Ideas of all those Qualities which they who make the Distinction, term Secondary?

Hyl. What! is it not an easy Matter to consider Extension and Motion by themselves, abstracted from all other sensible Qualities? Pray how do the Mathematicians treat of them?

Phil. I acknowledge, Hylas, it is not difficult to form general Propositions and Reasonings about those Qualities, without mentioning any other; and in this Sense to consider or treat of them abstractly. But how doth it follow that because I can pronounce the Word Motion by itself, I can form the Idea of it in my Mind exclusive of Body? Or because Theoremes may be made of Extension and Figures, without any mention of great or small, or any other sensible Mode or Quality; that therefore it is possible such an abstract Idea of Extension, without any particular Size or Figure, or sensible Quality, should be distinctly formed, and apprehended by the Mind? Mathematicians treat of Quantity, without regarding what other sensible Qualities it is attended with, as being altogether indifferent to their Demonstrations. But when laying aside the Words, they contemplate the bare Ideas, I believe you will find, they are not the pure abstracted Ideas of Extension.

Hyl. But what say you to pure Intellect? May not abstracted Ideas be framed by that Faculty?

Phil. Since I cannot frame abstract Ideas at all, it is plain, I cannot frame them by the help of pure Intellect, whatsoever Faculty you understand by those Words. Besides, not to inquire into the Nature of pure Intellect and its spiritual Objects, as Virtue, Reason, God, or the like; thus much seems manifest, that sensible Things are only to be perceived by Sense, or represented by the Imagination. Figures therefore and Extension being originally perceived by Sense, do not belong to pure Intellect. But for your farther Satisfaction, try if you can frame the Idea of any Figure, abstracted from all Particularities of Size, or even from other sensible Qualities.

Hyl. Let me think a little——I do not find that I can.

Phil. And can you think it possible, that should really exist in Nature, which implies a Repugnancy in its Conception?

Hyl. By no means.

Phil. Since therefore it is impossible even for the Mind to disunite the Ideas of Extension and Motion from all other sensible Qualities, doth it not follow, that where the one exist, there necessarily the other exist likewise?

Hyl. It should seem so.

Phil. Consequently the very same Arguments which you admitted, as conclusive against the Secondary Qualities, are without any farther Application of Force against the Primary too. Besides, if you will trust your Senses, is it not plain all sensible Qualities coexist, or to them, appear as being in the same Place? Do they ever represent a Motion, or Figure, as being divested of all other visible and tangible Qualities?

Hyl. You need say no more on this Head. I am free to own, if there be no secret Error or Oversight in our Proceedings hitherto, that all sensible Qualities are alike to be denied Existence without the Mind. But my Fear is, that I have been too liberal in my former Concessions, or overlooked some Fallacy or other. In short, I did not take time to think.

Phil. For that matter, Hylas, you may take what time you please in reviewing the Progress of our Inquiry. You are at liberty to recover any Slips you might have made, or offer whatever you have omitted, which makes for your first Opinion.

Hyl. One great Oversight I take to be this: That I did not sufficiently distinguish the Object from the Sensation. Now though this latter may not exist without the Mind, yet it will not thence follow that the former cannot.

Phil. What Object do you mean? the object of the Senses?

Hyl. The same.

Phil. It is then immediately perceived.

Hyl. Right.

Phil. Make me to understand the Difference between what is immediately perceived, and a Sensation.

Hyl. The Sensation I take to be an Act of the Mind perceiving; beside which, there is something perceived; and this I call the Object. For Example, there is Red and Yellow on that Tulip. But then the Act of perceiving those Colours is in me only, and not in the Tulip.

Phil. What Tulip do you speak of? Is it that which you see?

Hyl. The same.

Phil. And what do you see beside Colour, Figure, and Extension?

Hyl. Nothing.

Phil. What you would say then is, that the Red and Yellow are coexistent with the Extension; is it not?

Hyl. That is not all; I would say, They have a real Existence without the Mind, in some unthinking Substance.

Phil. That the Colours are really in the Tulip which I see, is manifest. Neither can it be denied, that this Tulip may exist independent of your Mind or mine; but that any immediate Object of the Senses, that is, any Idea, or Combination of Ideas, should exist in an unthinking Substance, or exterior to all Minds, is in itself an evident Contradiction. Nor can I imagine how this follows from what you said just now, to wit that the Red and Yellow were on the Tulip you saw, since you do not pretend to see that unthinking Substance.

Hyl. You have an artful way, Philonous, of diverting our Inquiry from the Subject. Phil. I see you have no mind to be pressed that way. To return then to your Distinction

between Sensation and Object; if I take you right, you distinguish in every Perception two Things, the one an Action of the Mind, the other not.

Hyl. True. Phil. And this Action cannot exist in, or belong to any unthinking thing; but whatever

beside is implied in a Perception, may.

Hyl. That is my Meaning. Phil. So that if there was a Perception without any Act of the Mind, it were possible

such a Perception should exist in an unthinking Substance.

Hyl. I grant it. But it is impossible there should be such a Perception. Phil. When is the Mind said to be active? Hyl. When it produces, puts an end to, or changes any thing. Phil. Can the Mind produce, discontinue, or change any thing but by an Act of the Will? Hyl. It cannot. Phil. The Mind therefore is to be accounted active in its Perceptions, so far forth as

Volition is included in them.

Hyl. It is. Phil. In plucking this Flower, I am active, because I do it by the Motion of my Hand,

which was consequent upon my Volition; so likewise in applying it to my Nose. But is either of these Smelling?

Hyl. No. Phil. I act too in drawing the Air through my Nose; because my Breathing so rather

than otherwise, is the Effect of my Volition. But neither can this be called Smelling: For if it were, I should smell every time I breathed in that manner.

Hyl. True. Phil. Smelling then is somewhat consequent to all this. Hyl. It is. Phil. But I do not find my Will concerned any farther. Whatever more there is, as that

I perceive such a particular Smell or any Smell at all, this is independent of my Will, and therein I am altogether passive. Do you find it otherwise with you, Hylas?

Hyl. No, the very same. Phil. Then as to Seeing, is it not in your Power to open your Eyes, or keep them shut;

to turn them this or that way?

Hyl. Without doubt. Phil. But doth it in like manner depend on your Will, that in looking on this Flower, you

perceive White rather than any other Colour? Or directing your open Eyes towards yonder Part of the Heaven, can you avoid seeing the Sun? Or is Light or Darkness the Effect of your Volition?

Hyl. No certainly.

Phil. You are then in these Respects altogether passive.

Hyl. I am.

Phil. Tell me now, whether Seeing consists in perceiving Light and Colours, or in opening and turning the Eyes?

Hyl. Without doubt, in the former.

Phil. Since therefore you are in the very Perception of Light and Colours altogether passive, what is become of that Action you were speaking of, as an Ingredient in every Sensation? And doth it not follow from your own Concessions, that the Perception of Light and Colours, including no Action in it, may exist in an unperceiving Substance? And is not this a plain Contradiction?

Hyl. I know not what to think of it.

Phil. Besides, since you distinguish the Active and Passive in every Perception, you must do it in that of Pain. But how is it possible that Pain, be it as little active as you please, should exist in an unperceiving Substance? In short, do but consider the Point, and then confess ingenuously, whether Light and Colours, Tastes, Sounds, &c. are not all equally Passions or Sensations in the Soul. You may indeed call them external Objects, and give them in Words what Subsistence you please. But examine your own Thoughts, and then tell me whether it be not as I say?

Hyl. I acknowledge, Philonous, that upon a fair Observation of what passes in my Mind, I can discover nothing else, but that I am a thinking Being, affected with Variety of Sensations; neither is it possible to conceive how a Sensation should exist in an unperceiving Substance. But then on the other hand, when I look on sensible Things in a different View, considering them as so many Modes and Qualities, I find it necessary to suppose a material Substratum, without which they cannot be conceived to exist.

Phil. Material Substratum call you it? Pray, by which of your Senses came you acquainted with that Being?

Hyl. It is not itself sensible; its Modes and Qualities only being perceived by the Senses. Phil. I presume then, it was by Reflexion and Reason you obtained the Idea of it. Hyl. I do not pretend to any proper positive Idea of it. However I conclude it exists,

because Qualities cannot be conceived to exist without a Support.

Phil. It seems then you have only a relative Notion of it, or that you conceive it not otherwise than by conceiving the Relation it bears to sensible Qualities.

Hyl. Right.

Phil. Be pleased therefore to let me know wherein that Relation consists. Hyl. Is it not sufficiently expressed in the Term Substratum, or Substance? Phil. If so, the Word Substratum should import, that it is spread under the sensible

Qualities or Accidents.

Hyl. True.

Phil. And consequently under Extension.

Hyl. I own it.

Phil. It is therefore somewhat in its own Nature intirely distinct from Extension.

Hyl. I tell you, Extension is only a Mode, and Matter is something that supports Modes. And is it not evident the Thing supported is different from the Thing supporting?

Phil. So that something distinct from, and exclusive of Extension, is supposed to be the Substratum of Extension.

Hyl. Just so. Phil. Answer me, Hylas. Can a Thing be spread without Extension? or is not the Idea

of Extension necessarily included in Spreading?

Hyl. It is. Phil. Whatsoever therefore you suppose spread under any thing, must have in itself an

Extension distinct from the Extension of that Thing under which it is spread.

Hyl. It must. Phil. Consequently every corporeal Substance being the Substratum of Extension, must

have in itself another Extension by which it is qualified to be a Substratum: And so on to Infinity. And I ask whether this be not absurd in itself, and repugnant to what you granted just now, to wit, that the Substratum was something distinct from, and exclusive of Extension.

Hyl. Aye but, Philonous, you take me wrong. I do not mean that Matter is spread in a gross literal Sense under Extension. The Word Substratum is used only to express in general the same thing with Substance.

Phil. Well then, let us examine the Relation implied in the Term Substance. Is it not that it stands under Accidents?

Hyl. The very same. Phil. But that one thing may stand under or support another, must it not be extended? Hyl. It must. Phil. Is not therefore this Supposition liable to the same Absurdity with the former? Hyl. You still take Things in a strict literal Sense: That is not fair, Philonous. Phil. I am not for imposing any Sense on your Words: You are at Liberty to explain

them as you please. Only I beseech you, make me understand something by them. You tell me, Matter supports or stands under Accidents. How! is it as your Legs support your Body?

Hyl. No; that is the literal Sense. Phil. Pray let me know any Sense, literal or not literal, that you understand it in.——

How long must I wait for an Answer, Hylas?

Hyl. I declare I know not what to say. I once thought I understood it well enough what was meant by Matter’s supporting Accidents. But now the more I think on it, the less can I comprehend it; in short, I find that I know nothing of it.

Phil. It seems then you have no Idea at all, neither relative nor positive of Matter; you know neither what it is in itself, nor what Relation it bears to Accidents.

Hyl. I acknowledge it. Phil. And yet you asserted, that you could not conceive how Qualities or Accidents

should really exist, without conceiving at the same time a material Support of them. Hyl. I did. Phil. That is to say, when you conceive the real Existence of Qualities, you do withal

conceive something which you cannot conceive.

Hyl. It was wrong I own. But still I fear there is some Fallacy or other. Pray what think you of this? It is just come into my Head, that the Ground of all our Mistake lies in your treating of each Quality by itself. Now, I grant that each Quality cannot singly subsist without the Mind. Colour cannot without Extension, neither can Figure without some other sensible Quality. But as the several Qualities united or blended together form intire sensible Things, nothing hinders why such Things may not be supposed to exist without the Mind.

Phil. Either, Hylas, you are jesting, or have a very bad Memory. Though indeed we went through all the Qualities by Name one after another; yet my Arguments, or rather your Concessions no where tended to prove, that the Secondary Qualities did not subsist each alone by itself; but that they were not at all without the Mind. Indeed in treating of Figure and Motion, we concluded they could not exist without the Mind, because it was impossible even in Thought to separate them from all Secondary Qualities, so as to conceive them existing by themselves. But then this was not the only Argument made use of upon that Occasion. But (to pass by all that hath been hitherto said, and reckon it for nothing, if you will have it so) I am content to put the whole upon this Issue. If you can conceive it possible for any Mixture or Combination of Qualities, or any sensible Object whatever, to exist without the Mind, then I will grant it actually to be so.

Hyl. If it comes to that, the Point will soon be decided. What more easy than to conceive a Tree or House existing by itself, independent of, and unperceived by any Mind whatsoever? I do at this present time conceive them existing after that manner.

Phil. How say you, Hylas, can you see a thing which is at the same time unseen?

Hyl. No, that were a Contradiction.

Phil. Is it not as great a Contradiction to talk of conceiving a thing which is unconceived? Hyl. It is. Phil. The Tree or House therefore which you think of, is conceived by you. Hyl. How should it be otherwise? Phil. And what is conceived, is surely in the Mind. Hyl. Without question, that which is conceived is in the Mind. Phil. How then came you to say, you conceived a House or Tree existing independent

and out of all Minds whatsoever?

Hyl. That was I own an Oversight; but stay, let me consider what led me to it.——It is a pleasant Mistake enough. As I was thinking of a Tree in a solitary Place, where no one was present to see it, methought that was to conceive a Tree as existing unperceived or unthought of, not considering that I myself conceived it all the while. But now I plainly see, that all I can do is to frame Ideas in my own Mind. I may indeed conceive in my own Thoughts the Idea of a Tree, or a House, or a Mountain, but that is all. And this is far from proving, that I can conceive them existing out of the Minds of all Spirits.

Phil. You acknowledge then that you cannot possibly conceive, how any one corporeal sensible Thing should exist otherwise than in a Mind.

Hyl. I do.

Phil. And yet you will earnestly contend for the Truth of that which you cannot so much as conceive.

Hyl. I profess I know not what to think, but still there are some Scruples remain with me. Is it not certain I see Things at a Distance? Do we not perceive the Stars and Moon, for Example, to be a great way off? Is not this, I say, manifest to the Senses?

Phil. Do you not in a Dream too perceive those or the like Objects?

Hyl. I do.

Phil. And have they not then the same Appearance of being distant?

Hyl. They have. Phil. But you do not thence conclude the Apparitions in a Dream to be without the

Mind?

Hyl. By no means. Phil. You ought not therefore to conclude that sensible Objects are without the Mind,

from their Appearance or Manner wherein they are perceived.

Hyl. I acknowledge it. But doth not my Sense deceive me in those Cases? Phil. By no means. The Idea or Thing which you immediately perceive, neither Sense

nor Reason inform you that it actually exists without the Mind. By Sense you only know that you are affected with such certain Sensations of Light and Colours, &c. And these you will not say are without the Mind.

Hyl. True: But beside all that, do you not think the Sight suggests something of Outness or Distance?

Phil. Upon approaching a distant Object, do the visible Size and Figure change perpetually, or do they appear the same at all Distances?

Hyl. They are in a continual Change. Phil. Sight therefore doth not suggest or any way inform you, that the visible Object you

immediately perceive, exists at a Distance *, or will be perceived when you advance further onward, there being a continued Series of visible Objects succeeding each other, during the whole time of your Approach.

Hyl. It doth not; but still I know, upon seeing an Object, what Object I shall perceive after having passed over a certain Distance: No matter whether it be exactly the same or no: There is still something of Distance suggested in the Case.

Phil. Good Hylas, do but reflect a little upon the Point, and then tell me whether there be any more in it than this. From the Ideas you actually perceive by Sight, you have by Experience learned to collect what other Ideas you will (according to the standing Order of Nature) be affected with, after such a certain Succession of Time and Motion.

Hyl. Upon the whole, I take it to be nothing else. Phil. Now is it not plain, that if we suppose a Man born blind was on a sudden made to

see, he could at first have no Experience of what may be suggested by Sight?

Hyl. It is. Phil. He would not then according to you have any Notion of Distance annexed to the

Things he saw; but would take them for a new Set of Sensations existing only in his Mind. Hyl. It is undeniable. Phil. But to make it still more plain: Is not Distance a Line turned endwise to the Eye? Hyl. It is. Phil. And can a Line so situated be perceived by Sight? Hyl. It cannot. Phil. Doth it not therefore follow that Distance is not properly and immediately perceived

by Sight?

Hyl. It should seem so. Phil. Again, is it your Opinion that Colours are at a Distance? Hyl. It must be acknowledged, they are only in the Mind.

* See the Essay towards a new Theory of Vision; And its Vindication.

Phil. But do not Colours appear to the Eye as coexisting in the same place with Extension and Figures?

Hyl. They do.

Phil. How can you then conclude from Sight, that Figures exist without, when you acknowledge Colours do not; the sensible Appearance being the very same with regard to both?

Hyl. I know not what to answer.

Phil. But allowing that Distance was truly and immediately perceived by the Mind, yet it would not thence follow it existed out of the Mind. For whatever is immediately perceived is an Idea: And can any Idea exist out of the Mind?

Hyl. To suppose that, were absurd: But inform me, Philonous, can we perceive or know nothing beside our Ideas?

Phil. As for the rational deducing of Causes from Effects, that is beside our Inquiry. And by the Senses you can best tell, whether you perceive any thing which is not immediately perceived. And I ask you, whether the Things immediately perceived, are other than your own Sensations or Ideas? You have indeed more than once, in the Course of this Conversation, declared yourself on those Points; but you seem by this last Question to have departed from what you then thought.

Hyl. To speak the Truth, Philonous, I think there are two Kinds of Objects, the one perceived immediately, which are likewise called Ideas; the other are real Things or external Objects perceived by the Mediation of Ideas, which are their Images and Representions. Now I own, Ideas do not exist without the Mind; but the latter sort of Objects do. I am sorry I did not think of this Distinction sooner; it would probably have cut short your Discourse.

Phil. Are those external Objects perceived by Sense, or by some other Faculty?

Hyl. They are perceived by Sense.

Phil. How! is there any thing perceived by Sense, which is not immediately perceived? Hyl. Yes, Philonous, in some sort there is. For Example, when I look on a Picture or

Statue of Julius Cæsar, I may be said after a manner to perceive him (though not immediately) by my Senses.

Phil. It seems then, you will have our Ideas, which alone are immediately perceived, to be Pictures of external Things: And that these also are perceived by Sense, inasmuch as they have a Conformity or Resemblance to our Ideas.

Hyl. That is my Meaning.

Phil. And in the same way that Julius Cæsar, in himself invisible, is nevertheless perceived by Sight; real Things in themselves imperceptible, are perceived by Sense.

Hyl. In the very same.

Phil. Tell me, Hylas, when you behold the Picture of Julius Cæsar, do you see with your Eyes more than some Colours and Figures with a certain Symmetry and Composition of the whole?

Hyl. Nothing else.

Phil. And would not a Man, who had never known any thing of Julius Cæsar, see as much?

Hyl. He would.

Phil. Consequently he hath his Sight, and the Use of it, in as perfect a Degree as you. Hyl. I agree with you.

Phil. Whence comes it then that your Thoughts are directed to the Roman Emperor, and his are not? This cannot proceed from the Sensations or Ideas of Sense by you then perceived; since you acknowledge you have no Advantage over him in that respect. It should seem therefore to proceed from Reason and Memory: should it not?

Hyl. It should.

Phil. Consequently it will not follow from that Instance, that any thing is perceived by Sense which is not immediately perceived. Though I grant we may in one Acceptation be said to perceive sensible Things mediately by Sense: That is, when from a frequently perceived Connexion, the immediate Perception of Ideas by one Sense suggests to the Mind others perhaps belonging to another Sense, which are wont to be connected with them. For instance, when I hear a Coach drive along the Streets, immediately I perceive only the Sound; but from the Experience I have had that such a Sound is connected with a Coach, I am said to hear the Coach. It is nevertheless evident, that in truth and strictness, nothing can be heard but Sound: And the Coach is not then properly perceived by Sense, but suggested from Experience. So likewise when we are said to see a red-hot Bar of Iron; the Solidity and Heat of the Iron are not the Objects of Sight, but suggested to the Imagination by the Colour and Figure, which are properly perceived by that Sense. In short, those Things alone are actually and strictly perceived by any Sense, which would have been perceived, in case that same Sense had then been first conferred on us. As for other Things, it is plain they are only suggested to the Mind by Experience grounded on former Perceptions. But to return to your Comparison of Cæsar’s Picture, it is plain, if you keep to that, you must hold the real Things or Archetypes of our Ideas are not perceived by Sense, but by some internal Faculty of the Soul, as Reason or Memory. I would therefore fain know, what Arguments you can draw from Reason for the Existence of what you call real Things or material Objects. Or whether you remember to have seen them formerly as they are in themselves? or if you have heard or read of any one that did.

Hyl. I see, Philonous, you are disposed to Rallery; but that will never convince me. Phil. My Aim is only to learn from you, the way to come at the Knowledge of material

Beings. Whatever we perceive, is perceived either immediately or mediately: By Sense, or by Reason and Reflexion. But as you have excluded Sense, pray shew me what Reason you have to believe their Existence; or what medium you can possibly make use of, to prove it either to mine or your own Understanding.

Hyl. To deal ingenuously, Philonous, now I consider the Point, I do not find I can give you any good Reason for it. But thus much seems pretty plain, that it is at least possible such Things may really exist. And as long as there is no Absurdity in supposing them, I am resolved to believe as I did, till you bring good Reasons to the contrary.

Phil. What! is it come to this, that you only believe the Existence of material Objects, and that your Belief is founded barely on the Possibility of its being true? Then you will have me bring Reasons against it: Though another would think it reasonable, the Proof should lie on him who holds the Affirmative. And after all, this very Point which you are now resolved to maintain without any Reason, is in effect what you have more than once during this Discourse seen good Reason to give up. But to pass over all this; if I understand you rightly, you say our Ideas do not exist without the Mind; but that they are Copies, Images, or Representations of certain Originals that do.

Hyl. You take me right.

Phil. They are then like external Things.

Hyl. They are.

Phil. Have those Things a stable and permanent Nature independent of our Senses; or are they in a perpetual Change, upon our producing any Motions in our Bodies, suspending, exerting, or altering our Faculties or Organs of Sense?

Hyl. Real Things, it is plain, have a fixed and real Nature, which remains the same notwithstanding any Change in our Senses, or in the Posture and Motion of our Bodies; which indeed may affect the Ideas in our Minds, but it were absurd to think they had the same Effect on Things existing without the Mind.

Phil. How then is it possible, that Things perpetually fleeting and variable as our Ideas, should be Copies or Images of any thing fixed and constant? Or in other Words, since all sensible Qualities, as Size, Figure, Colour, &c. that is, our Ideas are continally changing upon every Alteration in the Distance, Medium, or Instruments of Sensation; how can any determinate material Objects be properly represented or painted forth by several distinct Things, each of which is so different from and unlike the rest? Or if you say it resembles some one only of our Ideas, how shall we be able to distinguish the true Copy from all the false ones?

Hyl. I profess, Philonous, I am at a loss. I know not what to say to this.

Phil. But neither is this all. Which are material Objects in themselves, perceptible or imperceptible?

Hyl. Properly and immediately nothing can be perceived but Ideas. All material Things therefore are in themselves insensible, and to be perceived only by their Ideas.

Phil. Ideas then are sensible, and their Archetypes or Originals insensible.

Hyl. Right.

Phil. But how can that which is sensible be like that which is insensible? Can a real thing in itself invisible be like a Colour; or a real thing which is not audible, be like a Sound? In a word, can any thing be like a Sensation or Idea, but another Sensation or Idea?

Hyl. I must own, I think not.

Phil. Is it possible there should be any doubt in the Point? Do you not perfectly know your own Ideas?

Hyl. I know them perfectly; since what I do not perceive or know, can be no part of my Idea.

Phil. Consider therefore, and examine them, and then tell me if there be any thing in them which can exist without the Mind: or if you can conceive any thing like them existing without the Mind.

Hyl. Upon Inquiry, I find it is impossible for me to conceive or understand how any thing but an Idea can be like an Idea. And it is most evident, that no Idea can exist without the Mind.

Phil. You are therefore by your Principles forced to deny the Reality of sensible Things, since you made it consist in an absolute Existence exterior to the Mind. That is to say, you are a downright Sceptic. So I have gained my Point, which was to shew your Principles led to Scepticism.

Hyl. For the present I am, if not intirely convinced, at least silenced.

Phil. I would fain know what more you would require in order to a perfect Conviction. Have you not had the Liberty of explaining yourself all manner of ways? Were any little Slips in Discourse laid hold and insisted on? Or were you not allowed to retract or reinforce any thing you had offered, as best served your Purpose? Hath not every thing you could say been heard and examined with all the Fairness imaginable? In a word, have you not in every Point been convinced out of your own Mouth? And if you can at present discover any Flaw in any of your former Concessions, or think of any remaining Subterfuge, any new Distinction, Colour, or Comment whatsoever, why do you not produce it?

Hyl. A little Patience, Philonous. I am at present so amazed to see myself ensnared, and as it were imprisoned in the Labyrinths you have drawn me into, that on the sudden it cannot be expected I should find my way out. You must give me time to look about me, and recollect myself.

Phil. Hark; is not this the College-Bell?

Hyl. It rings for Prayers.

Phil. We will go in then if you please, and meet here again to Morrow Morning. In the meantime you may employ your Thoughts on this Morning’s Discourse, and try if you can find any Fallacy in it, or invent any new means to extricate yourself.

Hyl. Agreed.

THE SECOND

DIALOGUE.

Hylas. I beg your Pardon, Philonous, for not meeting you sooner. All this Morning my Head was so filled with our late Conversation, that I had not leisure to think of the Time of the Day, or indeed of any thing else.

Philonous. I am glad you were so intent upon it, in Hopes if there were any Mistakes in your Concessions, or Fallacies in my Reasonings from them, you will now discover them to me.

Hyl. I assure you, I have done nothing ever since I saw you, but search after Mistakes and Fallacies, and with that View have minutely examined the whole Series of Yesterday’s Discourse: but all in vain, for the Notions it led me into, upon Review appear still more clear and evident; and the more I consider them, the more irresistibly do they force my Assent.

Phil. And is not this, think you, a Sign that they are genuine, that they proceed from Nature, and are conformable to right Reason? Truth and Beauty are in this alike, that the strictest Survey sets them both off to Advantage. While the false Lustre of Error and Disguise cannot endure being reviewed, or too nearly inspected.

Hyl. I own there is a great deal in what you say. Nor can any one be more intirely satisfied of the Truth of those odd Consequences, so long as I have in View the Reasonings that lead to them. But when these are out of my Thoughts, there seems on the other hand something so satisfactory, so natural and intelligible in the modern way of explaining Things, that I profess I know not how to reject it.

Phil. I know not what way you mean.

Hyl. I mean the way of accounting for our Sensations or Ideas.

Phil. How is that?

Hyl. It is supposed the Soul makes her Residence in some part of the Brain, from which the Nerves take their rise, and are thence extended to all Parts of the Body: And that outward Objects by the different Impressions they make on the Organs of Sense, communicate certain vibrative Motions to the Nerves; and these being filled with Spirits, propagate them to the Brain or Seat of the Soul, which according to the various Impressions or Traces thereby made in the Brain, is variously affected with Ideas.

Phil. And call you this an Explication of the manner whereby we are affected with Ideas? Hyl. Why not, Philonous, have you any thing to object against it? Phil. I would first know whether I rightly understand your Hypothesis. You make certain

Traces in the Brain to be the Causes or Occasions of our Ideas. Pray tell me, whether by the Brain you mean any sensible Thing?

Hyl. What else think you I could mean?

Phil. Sensible Things are all immediately perceivable; and those Things which are immediately perceivable, are Ideas; and these exist only in the Mind. Thus much you have, if I mistake not, long since agreed to.

Hyl. I do not deny it.

Phil. The Brain therefore you speak of, being a sensible Thing, exists only in the Mind. Now, I would fain know whether you think it reasonable to suppose, that one Idea or Thing existing in the Mind, occasions all other Ideas. And if you think so, pray how do you account for the Origin of that Primary Idea or Brain itself?

Hyl. I do not explain the Origin of our Ideas by that Brain which is perceivable to Sense, this being itself only a Combination of sensible Ideas, but by another which I imagine.

Phil. But are not Things imagined as truly in the Mind as Things perceived?

Hyl. I must confess they are.

Phil. It comes therefore to the same thing; and you have been all this while accounting for Ideas, by certain Motions or Impressions in the Brain, that is, by some Alterations in an Idea, whether sensible or imaginable, it matters not.

Hyl. I begin to suspect my Hypothesis.

Phil. Beside Spirits, all that we know or conceive are our own Ideas. When therefore you say, all Ideas are occasioned by Impressions in the Brain, do you conceive this Brain or no? If you do, then you talk of Ideas imprinted in an Idea, causing that same Idea, which is absurd. If you do not conceive it, you talk unintelligibly, instead of forming a reasonable Hypothesis.

Hyl. I now clearly see it was a mere Dream. There is nothing in it.

Phil. You need not be much concerned at it: for after all, this way of explaining Things, as you called it, could never have satisfied any reasonable Man. What Connexion is there between a Motion in the Nerves, and the Sensations of Sound or Colour in the Mind? or how is it possible these should be the Effect of that?

Hyl. But I could never think it had so little in it, as now it seems to have.

Phil. Well then, are you at length satisfied that no sensible Things have a real Existence; and that you are in truth an arrant Sceptic?

Hyl. It is too plain to be denied.

Phil. Look! are not the Fields covered with a delightful Verdure? Is there not something in the Woods and Groves, in the Rivers and clear Springs that sooths, that delights, that transports the Soul? At the Prospect of the wide and deep Ocean, or some huge Mountain whose Top is lost in the Clouds, or of an old gloomy Forest, are not our Minds filled with a pleasing Horror? Even in Rocks and Deserts, is there not an agreeable Wildness? How sincere a Pleasure is it to behold the natural Beauties of the Earth! To preserve and renew our Relish for them, is not the Veil of Night alternately drawn over her Face, and doth she not change her Dress with the Seasons? How aptly are the Elements disposed? What Variety and Use in the meanest Productions of Nature? What Delicacy, what Beauty, what Contrivance in animal and vegetable Bodies? How exquisitely are all Things suited, as well to their particular Ends, as to constitute apposite Parts of the Whole! And while they mutually aid and support, do they not also set off and illustrate each other? Raise now your Thoughts from this Ball of Earth, to all those glorious Luminaries that adorn the high Arch of Heaven. The Motion and Situation of the Planets, are they not admirable for Use and Order? Were those (miscalled Erratique) Globes ever known to stray, in their repeated Journeys through the pathless Void? Do they not measure Areas round the Sun ever proportioned to the Times? So fixed, so immutable are the Laws by which the unseen Author of Nature actuates the Universe. How vivid and radiant is the Lustre of the fixed Stars! How magnificent and rich that negligent Profusion, with which they appear to be scattered throughout the whole Azure Vault! Yet if you take the Telescope, it brings into your Sight a new Host of Stars that escape the naked Eye. Here they seem contiguous and minute, but to a nearer View immense Orbs of Light at various Distances, far sunk in the Abyss of Space. Now you must call Imagination to your Aid. The feeble narrow Sense cannot descry innumerable Worlds revolving round the central Fires; and in those Worlds the Energy of an all-perfect Mind displayed in endless Forms. But neither Sense nor Imagination are big enough to comprehend the boundless Extent with all its glittering Furniture. Though the labouring Mind exert and strain each Power to its utmost reach, there still stands out ungrasped a Surplusage immeasurable. Yet all the vast Bodies that compose this mighty Frame, how distant and remote soever, are by some secret Mechanism, some divine Art and Force linked in a mutual Dependence and Intercourse with each other, even with this Earth, which was almost slipt from my Thoughts, and lost in the Croud of Worlds. Is not the whole System immense, beautiful, glorious beyond Expression and beyond Thought! What treatment then do those Philosophers deserve, who would deprive these noble and delightful Scenes of all Reality? How should those Principles be entertained, that lead us to think all the visible Beauty of the Creation a false imaginary Glare? To be plain, can you expect this Scepticism of yours will not be thought extravagantly absurd by all Men of Sense?

Hyl. Other Men may think as they please: But for your part you have nothing to reproach me with. My Comfort is, you are as much a Sceptic as I am.

Phil. There, Hylas, I must beg leave to differ from you.

Hyl. What! have you all along agreed to the Premises, and do you now deny the Conclusion, and leave me to maintain those Paradoxes by myself which you led me into? This surely is not fair.

Phil. I deny that I agreed with you in those Notions that led to Scepticism. You indeed said, the Reality of sensible Things consisted in an absolute Existence out of the Minds of Spirits, or distinct from their being perceived. And pursuant to this Notion of Reality, you are obliged to deny sensible Things any real Existence: That is, according to your own Definition, you profess yourself a Sceptic. But I neither said nor thought the Reality of sensible Things was to be defined after that manner. To me it is evident, for the Reasons you allow of, that sensible Things cannot exist otherwise than in a Mind or Spirit. Whence I conclude, not that they have no real Existence, but that seeing they depend not on my Thought, and have an Existence distinct from being perceived by me, there must be some other Mind wherein they exist. As sure therefore as the sensible World really exists, so sure is there an infinite omnipresent Spirit who contains and supports it.

Hyl. What! this is no more than I and all Christians hold; nay, and all others too who believe there is a God, and that he knows and comprehends all Things.

Phil. Ay, but here lies the Difference. Men commonly believe that all Things are known or perceived by God, because they believe the Being of a God, whereas I on the other side, immediately and necessarily conclude the Being of a God, because all sensible Things must be perceived by him.

Hyl. But so long as we all believe the same thing, what matter is it how we come by that Belief?

Phil. But neither do we agree in the same Opinion. For Philosophers, though they acknowledge all corporeal Beings to be perceived by God, yet they attribute to them an absolute Subsistence distinct from their being perceived by any Mind whatever, which I do not. Besides, is there no Difference between saying, There is a God, therefore he perceives all Things: and saying, Sensible Things do really exist: and if they really exist, they are necessarily perceived by an infinite Mind: therefore there is an infinite Mind, or God? This furnishes you with a direct and immediate Demonstration, from a most evident Principle, of the Being of a God. Divines and Philosophers had proved beyond all Controversy, from the Beauty and Usefulness of the several Parts of the Creation, that it was the Workmanship of God. But that setting aside all Help of Astronomy and natural Philosophy, all Contemplation of the Contrivance, Order, and Adjustment of Things, an infinite Mind should be necessarily inferred from the bare Existence of the sensible World, is an Advantage peculiar to them only who have made this easy Reflexion: That the sensible World is that which we perceive by our several Senses; and that nothing is perceived by the Senses beside Ideas; and that no Idea or Archetype of an Idea can exist otherwise than in a Mind. You may now, without any laborious search into the Sciences, without any Subtilty of Reason, or tedious Length of Discourse, oppose and baffle the most strenuous Advocate for Atheism. Those miserable Refuges, whether in an eternal Succession of unthinking Causes and Effects, or in a fortuitous Concourse of Atoms; those wild Imaginings of Vanini, Hobbes, and Spinosa; in a word the whole System of Atheism, is it not intirely overthrown by this single Reflexion on the Repugnancy included in supposing the Whole, or any Part, even the most rude and shapeless of the visible World, to exist without a Mind? Let any one of those Abettors of Impiety but look into his own Thoughts, and there try if he can conceive how so much as a Rock, a Desert, a Chaos, or confused Jumble of Atoms; how any thing at all, either sensible or imaginable, can exist independent of a Mind, and he need go no farther to be convinced of his Folly. Can any thing be fairer than to put a Dispute on such an Issue, and leave it to a Man himself to see if he can conceive, even in Thought, what he holds to be true in Fact, and from a notional to allow it a real Existence?

Hyl. It cannot be denied, there is something highly serviceable to Religion in what you advance. But do you not think it looks very like a Notion entertained by some eminent Moderns, of seeing all things in God?

Phil. I would gladly know that Opinion; pray explain it to me.

Hyl. They conceive that the Soul being immaterial, is incapable of being united with material Things, so as to perceive them in themselves, but that she perceives them by her Union with the Substance of God, which being spiritual is therefore purely intelligible, or capable of being the immediate Object of a Spirit’s Thought. Besides, the Divine Essence contains in it Perfections correspondent to each created Being; and which are for that Reason proper to exhibit or represent them to the Mind.

Phil. I do not understand how our Ideas, which are Things altogether passive and inert, can be the Essence, or any Part (or like any Part) of the Essence or Substance of God, who is an impassive, indivisible, purely active Being. Many more Difficulties and Objections there are, which occur at first View against this Hypothesis; but I shall only add that it is liable to all the Absurdities of the common Hypotheses, in making a created World exist otherwise than in the Mind of a Spirit. Beside all which it hath this peculiar to itself; that it makes that material World serve to no Purpose. And if it pass for a good Argument against other Hypotheses in the Sciences, that they suppose Nature or the Divine Wisdom to make something in vain, or do that by tedious round-about Methods, which might have been performed in a much more easy and compendious way, what shall we think of that Hypothesis which supposes the whole World made in vain?

Hyl. But what say you, are not you too of Opinion that we see all Things in God? If I mistake not, what you advance comes near it.

Phil. Few Men think, yet all will have Opinions. Hence Mens Opinions are superficial and confused. It is nothing strange that Tenets, which in themselves are ever so different, should nevertheless be confounded with each other by those who do not consider them attentively. I shall not therefore be surprised, if some Men imagine that I run into the Enthusiasm of Malbranche, though in truth I am very remote from it. He builds on the most abstract general Ideas, which I intirely disclaim. He asserts an absolute external World, which I deny. He maintains that we are deceived by our Senses, and know not the real Natures or the true Forms and Figures of extended Beings; of all which I hold the direct contrary. So that upon the whole there are no Principles more fundamentally opposite than his and mine. It must be owned I intirely agree with what the holy Scripture saith, That in God we live, and move, and have our Being. But that we see Things in his Essence after the manner above set forth, I am far from believing. Take here in brief my Meaning. It is evident that the Things I perceive are my own Ideas, and that no Idea can exist unless it be in a Mind. Nor is it less plain that these Ideas or Things by me perceived, either themselves or their Archetypes, exist independently of my Mind, since I know myself not to be their Author, it being out of my power to determine at pleasure, what particular Ideas I shall be affected with upon opening my Eyes or Ears. They must therefore exist in some other Mind, whose Will it is they should be exhibited to me. The Things, I say, immediately perceived, are Ideas or Sensations, call them which you will. But how can any Idea or Sensation exist in, or be produced by, any thing but a Mind or Spirit? This indeed is inconceivable; and to assert that which is inconceivable, is to talk Nonsense: Is it not?

Hyl. Without doubt.

Phil. But on the other hand, it is very conceivable that they should exist in, and be produced by, a Spirit; since this is no more than I daily experience in myself, inasmuch as I perceive numberless Ideas; and by an Act of my Will can form a great Variety of them, and raise them up in my Imagination: Though it must be confessed, these Creatures of the Fancy are not altogether so distinct, so strong, vivid, and permanent, as those perceived by my Senses, which latter are called Real Things. From all which I conclude, there is a Mind which affects me every Moment with all the sensible Impressions I perceive. And from the Variety, Order, and Manner of these, I conclude the Author of them to be wise, powerful, and good beyond comprehension. Mark it well; I do not say, I see Things by perceiving that which represents them in the intelligible Substance of God. This I do not understand; but I say, The Things by me perceived are known by the Understanding, and produced by the Will, of an infinite Spirit. And is not all this most plain and evident? Is there any more in it, than what a little Observation of our own Minds, and that which passes in them not only enableth us to conceive, but also obligeth us to acknowledge?

Hyl. I think I understand you very clearly; and own the Proof you give of a Deity seems no less evident, than it is surprising. But allowing that God is the Supreme and Universal Cause of all Things, yet may there not be still a Third Nature besides Spirits and Ideas? May we not admit a subordinate and limited Cause of our Ideas? In a word, may there not for all that be Matter?

Phil. How often must I inculcate the same thing? You allow the Things immediately perceived by Sense to exist no where without the Mind: But there is nothing perceived by Sense, which is not perceived immediately: therefore there is nothing sensible that exists without the Mind. The Matter therefore which you still insist on, is something intelligible, I suppose; something that may be discovered by Reason, and not by Sense.

Hyl. You are in the right.

Phil. Pray let me know what Reasoning your Belief of Matter is grounded on; and what this Matter is in your present Sense of it.

Hyl. I find myself affected with various Ideas, whereof I know I am not the Cause; neither are they the Cause of themselves, or of one another, or capable of subsisting by themselves, as being altogether inactive, fleeting, dependent Beings. They have therefore some Cause distinct from me and them: Of which I pretend to know no more, than that it is the Cause of my Ideas. And this thing, whatever it be, I call Matter.

Phil. Tell me, Hylas, hath every one a Liberty to change the current proper Signification annexed to a common Name in any Language? For Example, suppose a Traveller should tell you, that in a certain Country Men might pass unhurt through the Fire; and, upon explaining himself, you find he meant by the Word Fire that which others call Water: Or if he should assert that there are Trees that walk upon two Legs, meaning Men by the term Trees. Would you think this reasonable?

Hyl. No; I should think it very absurd. Common Custom is the Standard of Propriety in Language. And for any Man to affect speaking improperly, is to pervert the Use of Speech, and can never serve to a better purpose, than to protract and multiply Disputes where there is no Difference in Opinion.

Phil. And doth not Matter, in the common current Acceptation of the Word, signify an extended, solid, moveable, unthinking, inactive Substance?

Hyl. It doth.

Phil. And hath it not been made evident, that no such Substance can possibly exist? And though it should be allowed to exist, yet how can that which is inactive be a Cause; or that which is unthinking be a Cause of Thought? You may indeed, if you please, annex to the Word Matter a contrary Meaning to what is vulgarly received; and tell me you understand by it an unextended, thinking, active Being, which is the Cause of our Ideas. But what else is this, than to play with Words, and run into that very Fault you just now condemned with so much Reason? I do by no means find fault with your Reasoning, in that you collect a Cause from the Phenomena: But I deny that the Cause deducible by Reason can properly be termed Matter.

Hyl. There is indeed something in what you say. But I am afraid you do not thoroughly comprehend my Meaning. I would by no means be thought to deny that God or an Infinite Spirit is the Supreme Cause of all things. All I contend for, is, that subordinate to the Supreme Agent there is a Cause of a limited and inferior Nature, which concurs in the Production of our Ideas, not by any Act of Will or Spiritual Efficiency, but by that Kind of Action which belongs to Matter, viz. Motion.

Phil. I find, you are at every Turn relapsing into your old exploded Conceit, of a moveable and consequently an extended Substance existing without the Mind. What! Have you already forgot you were convinced, or are you willing I should repeat what has been said on that Head? In truth this is not fair Dealing in you, still to suppose the Being of that which you have so often acknowledged to have no Being. But not to insist farther on what has been so largely handled, I ask whether all your Ideas are not perfectly passive and inert, including nothing of Action in them?

Hyl. They are.

Phil. And are sensible Qualities any thing else but Ideas? Hyl. How often have I acknowledged that they are not? Phil. But is not Motion a sensible Quality? Hyl. It is. Phil. Consequently it is no Action. Hyl. I agree with you. And indeed it is very plain, that when I stir my Finger, it remains

passive; but my Will which produced the Motion, is active.

Phil. Now I desire to know in the first place, whether Motion being allowed to be no Action, you can conceive any Action besides Volition: And in the second place, whether to say something and conceive nothing be not to talk Nonsense: And lastly, whether having considered the Premises, you do not perceive that to suppose any efficient or active Cause of our Ideas, other than Spirit, is highly absurd and unreasonable?

Hyl. I give up the Point intirely. But though Matter may not be a Cause, yet what hinders its being an Instrument subservient to the Supreme Agent in the Production of our Ideas?

Phil. An Instrument, say you; pray what may be the Figure, Springs, Wheels, and Motions of that Instrument?

Hyl. Those I pretend to determine nothing of, both the Substance and its Qualities being intirely unknown to me.

Phil. What? You are then of Opinion, it is made up of unknown Parts, that it hath unknown Motions, and an unknown Shape.

Hyl. I do not believe that it hath any Figure or Motion at all, being already convinced, that no sensible Qualities can exist in an unperceiving Substance.

Phil. But what Notion is it possible to frame of an Instrument void of all sensible Qualities, even Extension itself?

Hyl. I do not pretend to have any Notion of it.

Phil. And what reason have you to think, this unknown, this inconceivable Somewhat doth exist? Is it that you imagine God cannot act as well without it, or that you find by Experience the Use of some such thing, when you form Ideas in your own Mind?

Hyl. You are always teizing me for Reasons of my Belief. Pray what Reasons have you not to believe it?

Phil. It is to me a sufficient Reason not to believe the Existence of any thing, if I see no Reason for believing it. But not to insist on Reasons for believing, you will not so much as let me know what it is you would have me believe, since you say you have no manner of Notion of it. After all, let me intreat you to consider whether it be like a Philosopher, or even like a Man of common Sense, to pretend to believe you know not what, and you know not why.

Hyl. Hold, Philonous. When I tell you Matter is an Instrument, I do not mean altogether Nothing. It is true, I know not the particular Kind of Instrument; but however I have some Notion of Instrument in general, which I apply to it.

Phil. But what if it should prove that there is something, even in the most general Notion of Instrument, as taken in a distinct Sense from Cause, which makes the Use of it inconsistent with the Divine Attributes?

Hyl. Make that appear, and I shall give up the Point.

Phil. What mean you by the general Nature or Notion of Instrument?

Hyl. That which is common to all particular Instruments, composeth the general Notion. Phil. Is it not common to all Instruments, that they are applied to the doing those things

only, which cannot be performed by the mere Act of our Wills? Thus for instance, I never use an Instrument to move my Finger, because it is done by a Volition. But I should use one, if I were to remove part of a Rock, or tear up a Tree by the Roots. Are you of the same Mind? Or can you shew any Example where an Instrument is made use of in producing an Effect immediately depending on the Will of the Agent?

Hyl. I own, I cannot.

Phil. How therefore can you suppose, that an All-perfect Spirit, on whose Will all things have an absolute and immediate Dependence, should need an Instrument in his Operations, or not needing it make use of it? Thus it seems to me that you are obliged to own the Use of a lifeless inactive Instrument, to be incompatible with the Infinite Perfection of God; that is, by your own Confession, to give up the Point.

Hyl. It doth not readily occur what I can answer you.

Phil. But methinks you should be ready to own the Truth, when it hath been fairly proved to you. We indeed, who are Beings of Finite Powers, are forced to make use of Instruments. And the Use of an Instrument sheweth the Agent to be limited by Rules of another’s Prescription, and that he cannot obtain his End, but in such a Way and by such Conditions. Whence it seems a clear Consequence, that the supreme unlimited Agent useth no Tool or Instrument at all. The Will of an Omnipotent Spirit is no sooner exerted than executed, without the Application of Means, which, if they are employed by inferior Agents, it is not upon account of any real Efficacy that is in them, or necessary Aptitude to produce any Effect, but merely in compliance with the Laws of Nature, or those conditions prescribed to them by the first Cause, who is Himself above all Limitation or Prescription whatsoever.

Hyl. I will no longer maintain that Matter is an Instrument. However, I would not be understood to give up its Existence neither; since, notwithstanding what hath been said, it may still be an Occasion.

Phil. How many Shapes is your Matter to take? Or how often must it be proved not to exist, before you are content to part with it? But to say no more of this (though by all the Laws of Disputation I may justly blame you for so frequently changing the Signification of the principal Term) I would fain know what you mean by affirming that Matter is an Occasion, having already denied it to be a Cause. And when you have shewn in what Sense you understand Occasion, pray in the next place be pleased to shew me what Reason induceth you to believe there is such an Occasion of our Ideas.

Hyl. As to the first Point: By Occasion I mean an inactive unthinking Being, at the Presence whereof God excites Ideas in our Minds.

Phil. And what may be the Nature of that inactive unthinking Being?

Hyl. I know nothing of its Nature.

Phil. Proceed then to the second Point, and assign some Reason why we should allow an Existence to this inactive, unthinking, unknown thing.

Hyl. When we see Ideas produced in our Minds after an orderly and constant manner, it is natural to think they have some fixed and regular Occasions, at the Presence of which they are excited.

Phil. You acknowledge then God alone to be the Cause of our Ideas, and that he causes them at the Presence of those Occasions.

Hyl. That is my Opinion.

Phil. Those Things which you say are present to God, without doubt He perceives. Hyl. Certainly; otherwise they could not be to Him an Occasion of acting. Phil. Not to insist now on your making Sense of this Hypothesis, or answering all the

puzzling Questions and Difficulties it is liable to: I only ask whether the Order and Regularity observable in the Series of our Ideas, or the Course of Nature, be not sufficiently accounted for by the Wisdom and Power of God; and whether it doth not derogate from those Attributes, to suppose He is influenced, directed, or put in mind, when and what He is to act, by any unthinking Substance. And lastly whether, in case I granted all you contend for, it would make any thing to your purpose, it not being easy to conceive how the external or absolute Existence of an unthinking Substance, distinct from its being perceived, can be inferred from my allowing that there are certain things perceived by the Mind of God, which are to Him the Occasion of producing Ideas in us.

Hyl. I am perfectly at a loss what to think, this Notion of Occasion seeming now altogether as groundless as the rest.

Phil. Do you not at length perceive, that in all these different Acceptations of Matter, you have been only supposing you know not what, for no manner of Reason, and to no kind of Use?

Hyl. I freely own my self less fond of my Notions, since they have been so accurately examined. But still, methinks I have some confused Perception that there is such a thing as Matter.

Phil. Either you perceive the Being of Matter immediately, or mediately. If immediately, pray inform me by which of the Senses you perceive it. If mediately, let me know by what Reasoning it is inferred from those Things which you perceive immediately. So much for the Perception. Then for the Matter it self, I ask whether it is Object, Substratum, Cause, Instrument, or Occasion? You have already pleaded for each of these, shifting your Notions, and making Matter to appear sometimes in one Shape, then in another. And what you have offered hath been disapproved and rejected by your self. If you have any thing new to advance, I would gladly hear it.

Hyl. I think I have already offered all I had to say on those Heads. I am at a loss what more to urge.

Phil. And yet you are loth to part with your old Prejudice. But to make you quit it more easily, I desire that, beside what has been hitherto suggested, you will farther consider whether, upon supposition that Matter exists, you can possibly conceive how you should be affected by it? Or supposing it did not exist, whether it be not evident you might for all that be affected with the same Ideas you now are, and consequently have the very same reasons to believe its Existence that you now can have?

Hyl. I acknowledge it is possible we might perceive all things just as we do now, though there was no Matter in the World; neither can I conceive, if there be Matter, how it should produce any Idea in our Minds. And I do farther grant, you have intirely satisfied me, that it is impossible there should be such a thing as Matter in any of the foregoing Acceptations. But still I cannot help supposing that there is Matter in some sense or other. What that is I do not indeed pretend to determine.

Phil. I do not expect you should define exactly the Nature of that unknown Being. Only be pleased to tell me, whether it is a Substance: And if so, whether you can suppose a Substance without Accidents; or in case you suppose it to have Accidents or Qualities, I desire you will let me know what those Qualities are, at least what is meant by Matter’s supporting them.

Hyl. We have already argued on those Points. I have no more to say to them. But to prevent any farther Questions, let me tell you, I at present understand by Matter neither Substance nor Accident, thinking nor extended Being, neither Cause, Instrument, nor Occasion, but something intirely unknown, distinct from all these.

Phil. It seems then you include in your present Notion of Matter, nothing but the general abstract Idea of Entity.

Hyl. Nothing else, save only that I superadd to this general Idea the Negation of all those particular Things, Qualities, or Ideas that I perceive, imagine, or in any wise apprehend.

Phil. Pray where do you suppose this unknown Matter to exist?

Hyl. Oh Philonous! now you think you have entangled me; for if I say it exists in Place, then you will infer that it exists in the Mind, since it is agreed, that Place or Extension exists only in the Mind: But I am not ashamed to own my Ignorance. I know not where it exists; only I am sure it exists not in Place. There is a negative Answer for you: And you must expect no other to all the Questions you put for the future about Matter.

Phil. Since you will not tell me where it exists, be pleased to inform me after what Manner you suppose it to exist, or what you mean by its Existence.

Hyl. It neither thinks nor acts, neither perceives, nor is perceived.

Phil. But what is there positive in your abstracted Notion of its Existence?

Hyl. Upon a nice Observation, I do not find I have any positive Notion or Meaning at all. I tell you again, I am not ashamed to own my Ignorance. I know not what is meant by its Existence, or how its exists.

Phil. Continue, good Hylas, to act the same ingenuous Part, and tell me sincerely whether you can frame a distinct Idea of Entity in general, prescinded from and exclusive of all thinking and corporal Beings, all particular things whatsoever.

Hyl. Hold, let me think a little——I profess, Philonous, I do not find that I can. At first Glance methought I had some dilute and airy Notion of pure Entity in Abstract; but upon closer Attention it hath quite vanished out of Sight. The more I think on it, the more am I confirmed in my prudent Resolution of giving none but negative Answers, and not pretending to the least Degree of any positive Knowledge or Conception of Matter, its Where, its How, its Entity, or any thing belonging to it.

Phil. When therefore you speak of the Existence of Matter, you have not any Notion in your Mind.

Hyl. None at all.

Phil. Pray tell me if the Case stands not thus: At first, from a Belief of Material Substance you would have it that the immediate Objects existed without the Mind; then that they are Archetypes; then Causes; next Instruments; then Occasions: Lastly, something in general, which being interpreted proves nothing. So Matter comes to nothing. What think you, Hylas, is not this a fair Summary of your whole Proceeding?

Hyl. Be that as it will, yet I still insist upon it, that our not being able to conceive a Thing, is no Argument against its Existence.

Phil. That from a Cause, Effect, Operation, Sign, or other Circumstance, there may reasonably be inferred the Existence of a Thing not immediately perceived, and that it were absurd for any Man to argue against the Existence of that Thing, from his having no direct and positive Notion of it, I freely own. But where there is nothing of all this; where neither Reason nor Revelation induces us to believe the Existence of a Thing; where we have not even a relative Notion of it; where an Abstraction is made from perceiving and being perceived, from Spirit and Idea: Lastly, where there is not so much as the most inadequate or faint Idea pretended to: I will not indeed thence conclude against the Reality of any Notion or Existence of any thing: But my Inference shall be, that you mean nothing at all: That you employ words to no manner of Purpose, without any Design or Signification whatsoever. And I leave it to you to consider how mere Jargon should be treated.

Hyl. To deal frankly with you, Philonous, your Arguments seem in themselves unanswerable, but they have not so great an Effect on me as to produce that intire Conviction, that hearty Acquiescence which attends Demonstration. I find myself still relapsing into an obscure Surmise of I know not what, Matter.

Phil. But are you not sensible, Hylas, that two Things must concur to take away all Scruple, and work a plenary Assent in the Mind? Let a visible Object be set in never so clear a Light, yet if there is any Imperfection in the Sight, or if the Eye is not directed towards it, it will not be distinctly seen. And though a Demonstration be never so well grounded and fairly proposed, yet if there is withal a Stain of Prejudice, or a wrong Bias on the Understanding, can it be expected on a sudden to perceive clearly and adhere firmly to the Truth? No, there is need of Time and Pains: The Attention must be awakened and detained by a frequent Repetition of the same Thing placed oft in the same, oft in different Lights. I have said it already, and find I must still repeat and inculcate, that it is an unaccountable Licence you take in pretending to maintain you know not what, for you know not what Reason, to you know not what Purpose. Can this be paralleled in any Art or Science, any Sect or Profession of Men? Or is there any thing so barefacedly groundless and unreasonable to be met with even in the lowest of common Conversation? But perhaps you will still say, Matter may exist, though at the same time you neither know what is meant by Matter, or by its Existence. This indeed is surprizing, and the more so because it is altogether voluntary, you not being led to it by any one Reason; for I challenge you to shew me that Thing in Nature which needs Matter to explain or account for it.

Hyl. The Reality of Things cannot be maintained without supposing the Existence of Matter. And is not this, think you, a good Reason why I should be earnest in its Defence?

Phil. The Reality of Things! What Things, sensible or intelligible?

Hyl. Sensible Things.

Phil. My Glove, for Example?

Hyl. That or any other thing perceived by the Senses.

Phil. But to fix on some particular thing; is it not a sufficient Evidence to me of the Existence of this Glove, that I see it, and feel it, and wear it? Or if this will not do, how is it possible I should be assured of the Reality of this Thing, which I actually see in this Place, by supposing that some unknown Thing which I never did or can see, exists after an unknown manner, in an unknown place, or in no place at all? How can the supposed Reality of that which is intangible, be a Proof that any thing tangible really exists? Or of that which is invisible, that any visible thing, or in general of any thing which is imperceptible, that a Perceptible exists? Do but explain this, and I shall think nothing too hard for you.

Hyl. Upon the whole, I am content to own the Existence of Matter is highly improbable; but the direct and absolute Impossibility of it does not appear to me.

Phil. But granting Matter to be possible, yet upon that account merely it can have no more Claim to Existence, than a Golden Mountain or a Centaur.

Hyl. I acknowledge it; but still you do not deny it is possible; and that which is possible, for ought you know, may actually exist.

Phil. I deny it to be possible; And have, if I mistake not, evidently proved from your own Concessions that it is not. In the common Sense of the Word Matter, is there any more implied, than an extended, solid, figured, moveable Substance existing without the Mind? And have not you acknowledged over and over, that you have seen evident Reason for denying the Possibility of such a Substance?

Hyl. True, but that is only one Sense of the Term Matter.

Phil. But is it not the only proper genuine received Sense? And if Matter in such a Sense be proved impossible, may it not be thought with good Grounds absolutely impossible? Else how could any thing be proved impossible? Or indeed how could there be any Proof at all one way or the other, to a Man who takes the Liberty to unsettle and change the common Signification of Words?

Hyl. I thought Philosophers might be allowed to speak more accurately than the Vulgar, and were not always confined to the common Acceptation of a Term.

Phil. But this now mentioned is the common received Sense among Philosophers themselves. But not to insist on that, have you not been allowed to take Matter in what Sense you pleased? And have you not used this Privilege in the utmost Extent, sometimes intirely changing, at others leaving out or putting into the Definition of it whatever for the present best served your Design, contrary to all the known Rules of Reason and Logick? And hath not this shifting unfair Method of yours spun out our Dispute to an unnecessary Length; Matter having been particularly examined, and by your own Confession refuted in each of those Senses? And can any more be required to prove the absolute Impossibility of a Thing, than the proving it impossible in every particular Sense, that either you or any one else understands it in?

Hyl. But I am not so thoroughly satisfied that you have proved the Impossibility of Matter in the last most obscure abstracted and indefinite Sense.

Phil. When is a thing shewn to be impossible?

Hyl. When a Repugnancy is demonstrated between the Ideas comprehended in its Definition.

Phil. But where there are no Ideas, there no Repugnancy can be demonstrated between Ideas.

Hyl. I agree with you.

Phil. Now in that which you call the obscure indefinite Sense of the Word Matter, it is plain, by your own Confession, there was included no Idea at all, no Sense except an unknown Sense, which is the same as none. You are not therefore to expect I should prove a Repugnancy between Ideas where there are no Ideas; or the Impossibility of Matter taken in an unknown Sense, that is no Sense at all. My business was only to shew, you meant nothing; and this you were brought to own. So that in all your various Senses, you have been shewed either to mean nothing at all, or if any thing, an Absurdity. And if this be not sufficient to prove the Impossibility of a Thing, I desire you will let be know what is.

Hyl. I acknowledge you have proved that Matter is impossible; nor do I see what more can be said in defence of it. But at the same time that I give up this, I suspect all my other Notions. For surely none could be more seemingly evident than this once was: And yet it now seems as false and absurd as ever it did true before. But I think we have discussed the Point sufficiently for the present. The remaining Part of the Day I would willingly spend, in running over in my Thoughts the several Heads of this Morning’s Conversation, and to Morrow shall be glad to meet you here again about the same time.

Phil. I will not fail to attend you.

THE THIRD

DIALOGUE.

Philonous. Tell me, Hylas, What are the Fruits of Yesterday’s Meditation? Hath it confirmed you in the same Mind you were in at parting? or have you since seen Cause to change your Opinion?

Hyl. Truly my Opinion is, that all our Opinions are alike vain and uncertain. What we approve to day, we condemn to morrow. We keep a Stir about Knowledge, and spend our Lives in the Pursuit of it, when, alas! we know nothing all the while: nor do I think it possible for us ever to know any thing in this Life. Our Faculties are too narrow and too few. Nature certainly never intended us for Speculation.

Phil. What! say you we can know nothing, Hylas?

Hyl. There is not a single thing in the World, whereof we can know the real Nature, or what it is in itself.

Phil. Will you tell me I do not really know what Fire or Water is?

Hyl. You may indeed know that Fire appears hot, and Water fluid: But this is no more than knowing what Sensations are produced in your own Mind, upon the Application of Fire and Water to your Organs of Sense. Their internal Constitution, their true and real Nature, you are utterly in the dark as to that.

Phil. Do I not know this to be a real Stone that I stand on, and that which I see before my Eyes to be a real Tree?

Hyl. Know? No, it is impossible you or any Man alive should know it. All you know, is, that you have such a certain Idea or Appearance in your own Mind. But what is this to the real Tree or Stone? I tell you, that Colour, Figure, and Hardness, which you perceive, are not the real Natures of those Things, or in the least like them. The same may be said of all other real Things or corporeal Substances which compose the World. They have none of them any thing in themselves, like those sensible Qualities by us perceived. We should not therefore pretend to affirm or know any thing of them, as they are in their own Nature.

Phil. But surely, Hylas, I can distinguish Gold, for Example, from Iron: And how could this be if I knew not what either truly was?

Hyl. Believe me, Philonous, you can only distinguish between your own Ideas. That Yellowness, that Weight, and other sensible Qualities, think you they are really in the Gold? They are only relative to the Senses, and have no absolute Existence in Nature. And in pretending to distinguish the Species of real Things, by the Appearances in your Mind, you may perhaps act as wisely as he that should conclude two Men were of a different Species, because their Clothes were not of the same Colour.

Phil. It seems then we are altogether put off with the Appearances of Things, and those false ones too. The very Meat I eat, and the Cloth I wear, have nothing in them like what I see and feel.

Hyl. Even so.

Phil. But is it not strange the whole World should be thus imposed on, and so foolish as to believe their Senses? And yet I know not how it is, but Men eat, and drink, and sleep, and perform all the Offices of Life as comfortably and conveniently, as if they really knew the Things they are conversant about.

Hyl. They do so: But you know ordinary Practice does not require a Nicety of speculative Knowledge. Hence the Vulgar retain their Mistakes, and for all that, make a Shift to bustle through the Affairs of Life. But Philosophers know better things.

Phil. You mean, they know that they know nothing.

Hyl. That is the very Top and Perfection of Humane Knowledge.

Phil. But are you all this while in earnest, Hylas; and are you seriously persuaded that you know nothing real in the World? Suppose you are going to write, would you not call for Pen, Ink, and Paper, like another Man; and do you not know what it is you call for?

Hyl. How often must I tell you, that I know not the real Nature of any one thing in the Universe? I may indeed upon Occasion make use of Pen, Ink, and Paper. But what any one of them is in its own true Nature, I declare positively I know not. And the same is true with regard to every other corporeal thing. And, what is more, we are not only ignorant of the true and real Nature of Things, but even of their Existence. It cannot be denied that we perceive such certain Appearances or Ideas; but it cannot be concluded from thence that Bodies really exist. Nay, now I think on it, I must agreeably to my former Concessions farther declare, that it is impossible any real corporeal Thing should exist in Nature.

Phil. You amaze me. Was ever any thing more wild and extravagant than the Notions you now maintain: And is it not evident you are led into all these Extravagances by the Belief of material Substance? This makes you dream of those unknown Natures in every thing. It is this occasions your distinguishing between the Reality and sensible Appearance of Things. It is to this you are indebted for being ignorant of what every Body else knows perfectly well. Nor is this all: You are not only ignorant of the true Nature of every Thing, but you know not whether any thing really exists, or whether there are any true Natures at all; forasmuch as you attribute to your material Beings an absolute or external Existence, wherein you suppose their Reality consists. And as you are forced in the end to acknowledge such an Existence means either a direct Repugnancy, or nothing at all, it follows that you are obliged to pull down your own Hypothesis of material Substance, and positively to deny the real Existence of any Part of the Universe. And so you are plunged into the deepest and most deplorable Scepticism that ever Man was. Tell me, Hylas, is it not as I say?

Hyl. I agree with you. Material Substance was no more than an Hypothesis, and a false and groundless one too. I will no longer spend my Breath in defence of it. But whatever Hypothesis you advance, or whatsoever Scheme of Things you introduce in its stead, I doubt not it will appear every whit as false: Let me but be allowed to question you upon it. This is, suffer me to serve you in your own kind, and I warrant it shall conduct you through as many Perplexities and Contradictions, to the very same State of Scepticism that I my self am in at present.

Phil. I assure you, Hylas, I do not pretend to frame any Hypothesis at all. I am of a vulgar Cast, simple enough to believe my Senses, and leave Things as I find them. To be plain, it is my Opinion, that the real Things are those very Things I see and feel, and perceive by my Senses. These I know, and finding they answer all the Necessities and Purposes of Life, have no reason to be solicitous about any other unknown Beings. A Piece of sensible Bread, for Instance, would stay my Stomach better than ten thousand times as much of that insensible, unintelligible, real Bread you speak of. It is likewise my Opinion, that Colours and other sensible Qualities are on the Objects. I cannot for my Life help thinking that Snow is white, and Fire hot. You indeed, who by Snow and Fire mean certain external, unperceived, unperceiving Substances, are in the right to deny Whiteness or Heat to be Affections inherent in them. But I, who understand by those Words the Things I see and feel, am obliged to think like other Folks. And as I am no Sceptic with regard to the Nature of Things, so neither am I as to their Existence. That a thing should be really perceived by my Senses, and at the same time not really exist, is to me a plain Contradiction; since I cannot prescind or abstract, even in Thought, the Existence of a sensible Thing from its being perceived. Wood, Stones, Fire, Water, Flesh, Iron, and the like Things, which I name and discourse of, are Things that I know. And I should not have known them, but that I perceived them by my Senses; and Things perceived by the Senses are immediately perceived; and Things immediately perceived are Ideas; and Ideas cannot exist without the Mind; their Existence therefore consists in being perceived; when therefore they are actually perceived, there can be no doubt of their Existence. Away then with all that Scepticism, all those ridiculous philosophical Doubts. What a Jest is it for a Philosopher to question the Existence of sensible Things, till he hath it proved to him from the Veracity of God: Or to pretend our Knowledge in this Point falls short of Intuition or Demonstration? I might as well doubt of my own Being, as of the Being of those Things I actually see and feel.

Hyl. Not so fast, Philonous: you say you cannot conceive how sensible Things should exist without the Mind. Do you not?

Phil. I do.

Hyl. Supposing you were annihilated, cannot you conceive it possible, that Things perceivable by Sense may still exist?

Phil. I can; but then it must be in another mind. When I deny sensible Things an Existence out of the Mind, I do not mean my Mind in particular, but all Minds. Now it is plain they have an Existence exterior to my Mind, since I find them by Experience to be independent of it. There is therefore some other Mind wherein they exist, during the Intervals between the Times of my perceiving them: As likewise they did before my Birth, and would do after my supposed Annihilation. And as the same is true, with regard to all other finite created Spirits; it necessarily follows, there is an Omnipresent Eternal Mind, which knows and comprehends all things, and exhibits them to our View in such a manner, and according to such Rules as He Himself hath ordained, and are by us termed the Laws of Nature.

Hyl. Answer me, Philonous. Are all our Ideas perfectly inert Beings? Or have they any Agency included in them?

Phil. They are altogether passive and inert.

Hyl. And is not God an Agent, a Being purely active?

Phil. I acknowledge it.

Hyl. No Idea therefore can be like unto, or represent the Nature of God.

Phil. It cannot.

Hyl. Since therefore you have no Idea of the Mind of God, how can you conceive it possible, that things should exist in his Mind? Or, if you can conceive the Mind of God, without having an Idea of it, why may not I be allowed to conceive the Existence of Matter, notwithstanding that I have no Idea of it?

Phil. As to your first Question; I own I have properly no Idea, either of God or any other Spirit; for these being active, cannot be represented by things perfectly inert, as our Ideas are. I do nevertheless know, that I who am a Spirit or thinking Substance, exist as certainly, as I know my Ideas exist. Farther, I know what I mean by the terms I and Myself ; and I know this immediately, or intuitively, though I do not perceive it as I perceive a Triangle, a Colour, or a Sound. The Mind, Spirit or Soul, is that indivisible unextended Thing, which thinks, acts, and perceives. I say indivisible, because unextended; and unextended, because extended, figured, moveable Things, are Ideas; and that which perceives Ideas, which thinks and wills, is plainly it self no Idea, nor like an Idea. Ideas are Things inactive, and perceived: And Spirits a sort of Beings altogether different from them. I do not therefore say my Soul is an Idea, or like an Idea. However, taking the Word Idea in a large Sense, my Soul can be said to furnish me with an Idea, that is, an Image, or Likeness of God, though indeed extremely inadequate. For all the Notion I have of God, is obtained by reflecting on my own Soul heightning its Powers, and removing its Imperfections. I have therefore, though not an inactive Idea, yet in my self some sort of an active thinking Image of the Deity. And though I perceive Him not by Sense, yet I have a Notion of Him, or know him by Reflexion and Reasoning. My own Mind and my own Ideas I have an immediate Knowledge of; and by the help of these, do mediately apprehend the Possibility of the Existence of other Spirits and Ideas. Farther, from my own Being, and from the Dependency I find in my self and my Ideas, I do by an Act of Reason, necessarily infer the Existence of a God, and all created Things in the Mind of God. So much for your first Question. For the second: I suppose by this time you can answer it your self. For you neither perceive Matter objectively, as you do an inactive Being or Idea, nor know it, as you do your self by a reflex Act: Neither do you mediately apprehend it by Similitude of the one or the other: Nor yet collect it by Reasoning from that which you know immediately. All which makes the Case of Matter widely different from that of the Deity.

Hyl. You say your own Soul supplies you with some sort of an Idea or Image of God. But at the same time you acknowledge you have, properly speaking, no Idea of your own Soul. You even affirm that Spirits are a sort of Beings altogether different from Ideas. Consequently that no Idea can be like a Spirit. We have therefore no Idea of any Spirit. You admit nevertheless that there is spiritual Substance, although you have no Idea of it; while you deny there can be such a thing as material Substance, because you have no Notion or Idea of it. Is this fair Dealing? To act consistently, you must either admit Matter or reject Spirit. What say you to this?

Phil. I say in the first place, that I do not deny the Existence of material Substance, merely because I have no Notion of it, but because the Notion of it is inconsistent, or in other words, because it is repugnant that there should be a Notion of it. Many things, for ought I know, may exist, whereof neither I nor any other Man hath or can have any Idea or Notion whatsoever. But then those things must be possible, that is, nothing inconsistent must be included in their Definition. I say secondly, that although we believe things to exist which we do not perceive; yet we may not believe that any particular thing exists, without some reason for such Belief: But I have no reason for believing the Existence of Matter. I have no immediate Intuition thereof: neither can I immediately from my Sensations, Ideas, Notions, Actions or Passions, infer an unthinking, unperceiving, inactive Substance, either by probable Deduction, or necessary Consequence. Whereas the Being of my self, that is, my own Soul, Mind or thinking Principle, I evidently know by Reflexion. You will forgive me if I repeat the same things in answer to the same Objections. In the very Notion or Definition of material Substance, there is included a manifest Repugnance and Inconsistency. But this cannot be said of the Notion of Spirit. That Ideas should exist in what doth not perceive, or be produced by what doth not act, is repugnant. But it is no Repugnancy to say, that a perceiving Thing should be the Subject of Ideas, or an active Thing the Cause of them. It is granted we have neither an immediate Evidence nor a demonstrative Knowledge of the Existence of other finite Spirits; but it will not thence follow that such Spirits are on a foot with material Substances: if to suppose the one to be inconsistent, and it be not inconsistent to suppose the other; if the one can be inferred by no Argument, and there is a Probability for the other; if we see Signs and Effects indicating distinct finite Agents like our selves, and see no Sign or Symptom whatever that leads to a rational Belief of Matter. I say lastly, that I have a Notion of Spirit, though I have not, strictly speaking, an Idea of it. I do not perceive it as an Idea or by Means of an Idea, but know it by Reflexion.

Hyl. Notwithstanding all you have said, to me it seems, that according to your own way of thinking, and in consequence of your own Principles, it should follow that you are only a System of floating Ideas, without any Substance to support them. Words are not to be used without a meaning. And as there is no more Meaning in spiritual Substance than in material Substance, the one is to be exploded as well as the other.

Phil. How often must I repeat, that I know or am conscious of my own Being; and that I my self am not my Ideas, but somewhat else, a thinking active Principle that perceives, knows, wills, and operates about Ideas. I know that I, one and the same self, perceive both Colours and Sounds: that a Colour cannot perceive a Sound, nor a Sound a Colour: That I am therefore one individual Principle, distinct from Colour and Sound; and, for the same reason, from all other sensible things and inert Ideas. But I am not in like manner conscious either of the Existence or Essence of Matter. On the contrary, I know that nothing inconsistent can exist, and that the Existence of Matter implies an Inconsistency. Farther, I know what I mean, when I affirm that there is a spiritual Substance or Support of Ideas, that is, That a Spirit knows and perceives Ideas. But I do not know what is meant, when it is said, that an unperceiving Substance hath inherent in it and supports either Ideas or the Archetypes of Ideas. There is therefore upon the whole no parity of case between Spirit and Matter.

Hyl. I own my self satisfied in this point. But do you in earnest think, the real Existence of sensible things consists in their being actually perceived? If so; How comes it that all Mankind distinguish between them? Ask the first Man you meet, and he shall tell you, to be perceived is one thing, and to exist is another.

Phil. I am content, Hylas, to appeal to the common Sense of the World for the Truth of my Notion. Ask the Gardiner, why he thinks yonder Cherry-Tree exists in the Garden, and he shall tell you, because he sees and feels it; in a word, because he perceives it by his Senses. Ask him, why he thinks an Orange-Tree not to be there, and he shall tell you, because he does not perceive it. What he perceives by Sense, that he terms a real Being, and saith it is, or exists; but that which is not perceivable, the same, he saith, hath no Being.

Hyl. Yes, Philonous, I grant the Existence of a sensible thing consists in being perceivable, but not in being actually perceived.

Phil. And what is perceivable but an Idea? And can an Idea exist without being actually perceived? These are Points long since agreed between us.

Hyl. But be your opinion never so true, yet surely you will not deny it is shocking, and contrary to the common Sense of Men. Ask the Fellow, whether yonder Tree hath an Existence out of his Mind: What Answer think you he would make?

Phil. The same that I would my self, to wit, That it doth exist out of his Mind. But then to a Christian it cannot surely be shocking to say, The real Tree existing without his Mind is truly known and comprehended by (that is, exists in) the infinite Mind of God. Probably he may not at first glance be aware of the direct and immediate Proof there is of this, inasmuch as the very Being of a Tree, or any other sensible Thing, implies a Mind wherein it is. But the Point it self he cannot deny. The Question between the Materialists and me is not, whether Things have a real Existence out of the Mind of this or that Person, but whether they have an absolute Existence, distinct from being perceived by God, and exterior to all Minds. This indeed some Heathens and Philosophers have affirmed, but whoever entertains Notions of the Deity suitable to the Holy Scriptures, will be of another Opinion.

Hyl. But according to your Notions, what Difference is there between real Things, and Chimeras formed by the Imagination, or the Visions of a Dream, since they are all equally in the Mind?

Phil. The Ideas formed by the Imagination are faint and indistinct; they have besides an intire Dependence on the Will. But the Ideas perceived by Sense, that is, real Things, are more vivid and clear, and being imprinted on the Mind by a Spirit distinct from us, have not the like Dependence on our Will. There is therefore no Danger of confounding these with the foregoing: and there is as little of confounding them with the Visions of a Dream, which are dim, irregular, and confused. And though they should happen to be never so lively and natural, yet by their not being connected, and of a piece with the preceding and subsequent Transactions of our Lives, they might easily be distinguished from Realities. In short, by whatever Method you distinguish Things from Chimeras on your own Scheme, the same, it is evident, will hold also upon mine. For it must be, I presume, by some perceived Difference, and I am not for depriving you of any one thing that you perceive.

Hyl. But still, Philonous, you hold, there is nothing in the World but Spirits and Ideas. And this, you must needs acknowledge, sounds very odly.

Phil. I own the word Idea, not being commonly used for Thing, sounds something out of the way. My Reason for using it was, because a necessary Relation to the Mind is understood to be implied by that Term; and it is now commonly used by Philosophers, to denote the immediate Objects of the Understanding. But however odly the Proposition may sound in Words, yet it includes nothing so very strange or shocking in its Sense, which in effect amounts to no more than this, to wit, that there are only Things perceiving, and Things perceived; or that every unthinking Being is necessarily, and from the very Nature of its Existence, perceived by some Mind; if not by a finite created Mind, yet certainly by the infinite Mind of God, in whom we live, and move, and have our Being. Is this as strange as to say, The sensible Qualities are not on the Objects: Or, That we cannot be sure of the Existence of Things, or know any thing of their real Natures, though we both see and feel them, and perceive them by all our Senses?

Hyl. And in Consequence of this, must we not think there are no such Things as Physical or Corporeal Causes; but that a Spirit is the immediate Cause of all the Phænomena in Nature? Can there by any thing more extravagant than this?

Phil. Yes, it is infinitely more extravagant to say, A thing which is inert, operates on the Mind, and which is unperceiving, is the Cause of our Perceptions. Besides, that which to you, I know not for what Reason, seems so extravagant, is no more than the Holy Scriptures assert in a hundred Places. In them God is represented as the sole and immediate Author of all those Effects, which some Heathens and Philosophers are wont to ascribe to Nature, Matter, Fate, or the like unthinking Principle. This is so much the constant Language of Scripture, that it were needless to confirm it by Citations.

Hyl. You are not aware, Philonous, that in making God the immediate Author of all the Motions in Nature, you make him the Author of Murder, Sacrilege, Adultery, and the like heinous Sins.

Phil. In Answer to that, I observe first, that the Imputation of Guilt is the same, whether a Person commits an Action with or without an Instrument. In case therefore you suppose God to act by the Mediation of an Instrument, or Occasion, called Matter, you as truly make Him the Author of Sin as I, who think Him the immediate Agent in all those Operations vulgarly ascribed to Nature. I farther observe, that Sin or moral Turpitude doth not consist in the outward Physical Action or Motion, but in the internal Deviation of the Will from the Laws of Reason and Religion. This is plain, in that the killing an Enemy in a Battle, or putting a Criminal legally to Death, is not thought sinful, though the outward Act be the very same with that in the Case of Murder. Since therefore Sin doth not consist in the Physical Action, the making God an immediate Cause of all such Actions, is not making him the Author of Sin. Lastly, I have no where said that God is the only Agent who produces all the Motions in Bodies. It is true, I have denied there are any other Agents besides Spirits: But this is very consistent with allowing to Thinking Rational Beings, in the Production of Motions, the Use of limited Powers, ultimately indeed derived from God, but immediately under the Direction of their own Wills, which is sufficient to intitle them to all the Guilt of their Actions.

Hyl. But the denying Matter, Philonous, or corporeal Substance; there is the Point. You can never persuade me that this is not repugnant to the universal Sense of Mankind. Were our Dispute to be determined by most Voices, I am confident you would give up the Point, without gathering the Votes.

Phil. I wish both our Opinions were fairly stated and submitted to the Judgment of Men who had plain common Sense, without the Prejudices of a learned Education. Let me be represented as one who trusts his Senses, who thinks he knows the Things he sees and feels, and entertains no Doubts, of their Existence; and you fairly set forth with all your Doubts, your Paradoxes, and your Scepticism about you, and I shall willingly acquiesce in the Determination of any indifferent Person. That there is no Substance wherein Ideas can exist beside Spirit, is to me evident. And that the Objects immediately perceived are Ideas, is on all Hands agreed. And that sensible Qualities are Objects immediately perceived, no one can deny. It is evident there can be no Substratum of those Qualities but Spirit, in which they exist, not by way of Mode or Property, but as a thing perceived in that which perceives it. I deny therefore that there is any unthinking Substratum of the Objects of Sense, and in that Acceptation that there is any material Substance. But if by material Substance is meant only sensible Body, that which is seen and felt, (and the unphilosophical Part of the World, I dare say, mean no more) then I am more certain of Matter’s Existence than you, or any other Philosopher, pretend to be. If there be any thing which makes the Generality of Mankind averse from the Notions I espouse, it is a Misapprehension that I deny the Reality of sensible Things: But as it is you who are guilty of that and not I, it follows that in truth their Aversion is against your Notions, and not mine. I do therefore assert that I am as certain as of my own Being, that there are Bodies or corporeal Substances, (meaning the Things I perceive by my Senses) and that granting this, the Bulk of Mankind will take no Thought about, nor think themselves at all concerned in the Fate of those unknown Natures, and Philosophical Quiddities, which some Men are so fond of.

Hyl. What say you to this? Since, according to you, Men judge of the Reality of Things by their Senses, how can a Man be mistaken in thinking the Moon a plain lucid Surface, about a Foot in Diameter; or a square Tower, seen at a distance, round; or an Oar, with one End in the Water, crooked?

Phil. He is not mistaken with regard to the Ideas he actually perceives; but in the Inferences he makes from his present Perceptions. Thus in the Case of the Oar, what he immediately perceives by Sight is certainly crooked; and so far he is in the right. But if he thence conclude, that on taking the Oar out of the Water he shall perceive the same Crookedness; or that it would affect his Touch, as crooked things are wont to do: In that he is mistaken. In like manner, if he shall conclude from what he perceives in one Station, that in case he advances toward the Moon or Tower, he should still be affected with the like Ideas, he is mistaken. But his Mistake lies not in what he perceives immediately and at present, (it being a manifest Contradiction to suppose he should err in respect of that) but in the wrong Judgment he makes concerning the Ideas he apprehends to be connected with those immediately perceived: Or concerning the Ideas that, from what he perceives at present, he imagines would be perceived in other Circumstances. The Case is the same with regard to the Copernican system. We do not here perceive any Motion of the Earth: But it were erroneous thence to conclude, that in case we were placed at as great a Distance from that, as we are now from the other Planets, we should not then perceive its Motion.

Hyl. I understand you; and must needs own you say things plausible enough: But give me leave to put you in mind of one thing. Pray, Philonous, were you not formerly as positive that Matter existed, as you are now that it does not?

Phil. I was. But here lies the Difference. Before, my Positiveness was founded without Examination, upon Prejudice; but now, after Inquiry, upon Evidence.

Hyl. After all, it seems our Dispute is rather about Words than Things. We agree in the Thing, but differ in the Name. That we are affected with Ideas from without is evident; and it is no less evident, that there must be (I will not say Archetypes, but) Powers without the Mind, corresponding to those Ideas. And as these Powers cannot subsist by themselves, there is some Subject of them necessarily to be admitted, which I call Matter, and you call Spirit. This is all the Difference.

Phil. Pray, Hylas, is that powerful Being, or Subject of Powers, extended?

Hyl. It hath not Extension; but it hath the Power to raise in you the Idea of Extension. Phil. It is therefore itself unextended. Hyl. I grant it. Phil. Is it not also active? Hyl. Without doubt: Otherwise, how could we attribute Powers to it? Phil. Now let me ask you Two Questions: First, Whether it be agreeable to the Usage

either of Philosophers or others, to give the Name Matter to an unextended active Being?

And Secondly, Whether it be not ridiculously absurd to misapply Names contrary to the common Use of Language?

Hyl. Well then, let it not be called Matter, since you will have it so, but some Third Nature distinct from Matter and Spirit. For what reason is there why you should call it Spirit? Does not the Notion of Spirit imply, that it is thinking as well as active and unextended?

Phil. My Reason is this: because I have a mind to have some Notion or Meaning in what I say: but I have no Notion of any Action distinct from Volition, neither can I conceive Volition to be any where but in a Spirit: therefore when I speak of an active Being, I am obliged to mean a Spirit. Beside, what can be plainer than that a thing which hath no Ideas in itself, cannot impart them to me; and if it hath Ideas, surely it must be a Spirit. To make you comprehend the Point still more clearly if it be possible: I assert as well as you, that since we are affected from without, we must allow Powers to be without in a Being distinct from ourselves. So far we are agreed. But then we differ as to the Kind of this powerful Being. I will have it to be Spirit, you Matter, or I know not what (I may add too, you know not what) Third Nature. Thus I prove it to be a Spirit. From the Effects I see produced, I conclude there are Actions; and because Actions, Volitions; and because there are Volitions, there must be a Will. Again, the Things I perceive must have an Existence, they or their Archetypes, out of my Mind: But being Ideas, neither they nor their Archetypes can exist otherwise than in an Understanding: There is therefore an Understanding. But Will and Understanding constitute in the strictest Sense a Mind or Spirit. The powerful Cause therefore of my Ideas, is in strict Propriety of Speech a Spirit.

Hyl. And now I warrant you think you have made the Point very clear, little suspecting that what you advance leads directly to a Contradiction. Is it not an Absurdity to imagine any Imperfection in God?

Phil. Without doubt.

Hyl. To suffer Pain is an Imperfection.

Phil. It is.

Hyl. Are we not sometimes affected with Pain and Uneasiness by some other Being? Phil. We are. Hyl. And have you not said that Being is a Spirit, and is not that Spirit God? Phil. I grant it. Hyl. But you have asserted, that whatever Ideas we perceive from without, are in the

Mind which affects us. The Ideas therefore of Pain and Uneasiness are in God; or in other words, God suffers Pain: That is to say, there is an Imperfection in the Divine Nature, which you acknowledged was absurd. So you are caught in a plain Contradiction.

Phil. That God knows or understands all things, and that He knows among other things what Pain is, even every sort of painful Sensation, and what it is for His Creatures to suffer Pain, I make no question. But that God, though He knows and sometimes causes painful Sensations in us, can Himself suffer Pain, I positively deny. We who are limited and dependent Spirits, are liable to Impressions of Sense, the Effects of an external Agent, which being produced against our Wills, are sometimes painful and uneasy. But God, whom no external Being can affect, who perceives nothing by Sense as we do, whose Will is absolute and independent, causing all things, and liable to be thwarted or resisted by nothing; it is evident, such a Being as this can suffer nothing, nor be affected with any painful Sensation, or indeed any Sensation at all. We are chained to a Body, that is to say, our Perceptions are connected with corporeal Motions. By the Law of our Nature we are affected upon every Alteration in the nervous Parts of our sensible Body: Which sensible Body rightly considered, is nothing but a Complexion of such Qualities or Ideas, as have no Existence distinct from being perceived by a Mind: So that this Connexion of Sensations with corporeal Motions, means no more than a Correspondence in the Order of Nature between two Sets of Ideas, or Things immediately perceivable. But God is a pure Spirit, disengaged from all such Sympathy or natural Ties. No corporeal Motions are attended with the Sensations of Pain or Pleasure in his Mind. To know every thing knowable is certainly a Perfection; but to endure, or suffer, or feel any thing by Sense, is an Imperfection. The former, I say, agrees to God, but not the latter. God knows or hath Ideas; but His Ideas are not convey’d to Him by Sense, as ours are. Your not Distinguishing where there is so manifest a Difference, makes you fancy you see an Absurdity where there is none.

Hyl. But all this while you have not considered, that the Quantity of Matter hath been demonstrated to be proportional to the Gravity of Bodies. And what can withstand Demonstration?

Phil. Let me see how you demonstrate that Point.

Hyl. I lay it down for a Principle, that the Moments or Quantities of Motion in Bodies, are in a direct compounded Reason of the Velocities and Quantities of Matter contained in them. Hence, where the Velocities are equal, it follows, the Moments are directly as the Quantity of Matter in each. But it is found by Experience, that all Bodies (bating the small Inequalities, arising from the Resistance of the Air) descend with an equal Velocity; the Motion therefore of descending Bodies, and consequently their Gravity, which is the Cause or Principle of that Motion, is proportional to the Quantity of Matter: which was to be demonstrated.

Phil. You lay it down as a self-evident Principle, that the Quantity of Motion in any Body, is proportional to the Velocity and Matter taken together: And this is made use of to prove a Proposition, from whence the Existence of Matter is inferred. Pray is not this arguing in a Circle?

Hyl. In the Premise I only mean, that the Motion is proportional to the Velocity, jointly with the Extension and Solidity.

Phil. But allowing this to be true, yet it will not thence follow, that Gravity is proportional to Matter, in your Philosophic Sense of the Word; except you take it for granted, that unknown Substratum, or whatever else you call it, is proportional to those sensible Qualities; which to suppose, is plainly begging the Question. That there is Magnitude and Solidity, or Resistance, perceived by Sense, I readily grant; as likewise that Gravity may be proportional to those Qualities, I will not dispute. But that either these Qualities as perceived by us, or the Powers producing them do exist in a material Substratum; this is what I deny, and you indeed affirm, but notwithstanding your Demonstration, have not yet proved.

Hyl. I shall insist no longer on that Point. Do you think however, you shall persuade me the natural Philosophers have been dreaming all this while; pray what becomes of all their Hypotheses and Explications of the Phænomena, which suppose the Existence of Matter?

Phil. What mean you, Hylas, by the Phænomena?

Hyl. I mean the Appearances which I perceive by my Senses.

Phil. And the Appearances perceived by Sense, are they not Ideas?

Hyl. I have told you so a hundred times.

Phil. Therefore, to explain the Phænomena, is to shew how we come to be affected with Ideas, in that Manner and Order wherein they are imprinted on our Senses. Is it not?

Hyl. It is.

Phil. Now if you can prove, that any Philosopher hath explained the Production of any one Idea in our Minds by the Help of Matter, I shall for ever acquiesce, and look on all that hath been said against it as nothing: But if you cannot, it is vain to urge the Explication of Phænomena. That a Being endowed with Knowledge and Will, should produce or exhibit Ideas, is easily understood. But that a Being which is utterly destitute of these Faculties should be able to produce Ideas, or in any sort to affect an Intelligence, this I can never understand. This I say, though we had some positive Conception of Matter, though we knew its Qualities, and could comprehend its Existence, would yet be so far from explaining things, that it is it self the most inexplicable thing in the World. And yet for all this, it will not follow, that Philosophers have been doing nothing; for by observing and reasoning upon the Connexion of Ideas, they discover the Laws and Methods of Nature, which is a part of Knowledge both useful and entertaining.

Hyl. After all, can it be supposed God would deceive all Mankind? Do you imagine, He would have induced the whole World to believe the Being of Matter, if there was no such thing?

Phil. That every epidemical Opinion arising from Prejudice, or Passion, or Thoughtlessness, may be imputed to God, as the Author of it, I believe you will not affirm. Whatsoever Opinion we father on Him, it must be either because He has discovered it to use by supernatural Revelation, or because it is so evident to our natural Faculties, which were framed and given us by God, that it is impossible we should withhold our Assent from it. But where is the Revelation? or where is the Evidence that extorts the Belief of Matter? Nay, how does it appear, that Matter taken for something distinct from what we perceive by our Senses, is thought to exist by all Mankind, or indeed by any except a few Philosophers, who do not know what they would be at? Your Question supposes these Points are clear; and when you have cleared them, I shall think my self obliged to give you another Answer. In the mean time let it suffice that I tell you, I do not suppose God has deceived Mankind at all.

Hyl. But the Novelty, Philonous, the Novelty! There lies the Danger. New Notions should always be discountenanced; they unsettle Men’s Minds, and no body knows where they will end.

Phil. Why the rejecting a Notion that hath no Foundation either in Sense or in Reason, or in Divine Authority, should be thought to unsettle the Belief of such Opinions as are grounded on all or any of these, I cannot imagine. That Innovations in Government and Religion, are dangerous, and ought to be discountenanced, I freely own. But is there the like Reason why they should be discouraged in Philosophy? The making any thing known which was unknown before, is an Innovation in Knowledge: And if all such Innovations had been forbidden, Men would have made a notable Progress in the Arts and Sciences. But it is none of my business to plead for Novelties and Paradoxes. That the Qualities we perceive, are not on the Objects: That we must not believe our Senses: That we know nothing of the real Nature of Things, and can never be assured even of their Existence: That real Colours and Sounds are nothing but certain unknown Figures and Motions: That Motions are in themselves neither swift nor slow: That there are in Bodies absolute Extensions, without any particular Magnitude or Figure: That a Thing stupid, thoughtless and inactive, operates on a Spirit: That the least Particle of a Body, contains innumerable extended Parts. These are the Novelties, these are the strange Notions which shock the genuine uncorrupted Judgment of all Mankind; and being once admitted, embarrass the Mind with endless Doubts and Difficulties. And it is against these and the like Innovations, I endeavour to vindicate common Sense. It is true, in doing this, I may perhaps be obliged to use some Ambages, and ways of Speech not common. But if my Notions are once thorowly understood, that which is most singular in them, will in effect be found to amount to no more than this: That it is absolutely impossible, and a plain Contradiction to suppose, any unthinking Being should exist without being perceived by a Mind. And if this Notion be singular, it is a shame it should be so at this time of day, and in a Christian Country.

Hyl. As for the Difficulties other Opinions may be liable to, those are out of the Question. It is your Business to defend your own Opinion. Can any thing be plainer, than that you are for changing all things into Ideas? You, I say, who are not ashamed to charge me with Scepticism. This is so plain, there is no denying it.

Phil. You mistake me. I am not for changing Things into Ideas, but rather Ideas into Things; since those immediate Objects of Perception, which according to you, are only Appearances of Things, I take to be the real Things themselves.

Hyl. Things! you may pretend what you please; but it is certain, you leave us nothing but the empty Forms of Things, the Outside only which strikes the Senses.

Phil. What you call the empty Forms and Outside of Things, seems to me the very Things themselves. Nor are they empty or incomplete otherwise, than upon your Supposition, that Matter is an essential Part of all corporeal Things. We both therefore agree in this, that we perceive only sensible Forms: But herein we differ, you will have them to be empty Appearances, I real Beings. In short you do not trust your Senses, I do.

Hyl. You say you believe your Senses; and seem to applaud your self that in this you agree with the Vulgar. According to you therefore, the true Nature of a Thing is discovered by the Senses. If so, whence comes that Disagreement? Why is not the same Figure, and other sensible Qualities, perceived all manner of Ways? and why should we use a Microscope, the better to discover the true Nature of a Body, if it were discoverable to the naked Eye?

Phil. Strictly speaking, Hylas, we do not see the same Object that we feel; neither is the same Object perceived by the Microscope, which was by the naked Eye. But in case every Variation was thought sufficient to constitute a new Kind or Individual, the endless Number or Confusion of Names would render Language impracticable. Therefore to avoid this as well as other Inconveniences which are obvious upon a little Thought, Men combine together several Ideas, apprehended by divers Senses, or by the same Sense at different times, or in different Circumstances, but observed however to have some Connexion in Nature, either with respect to Coexistence or Succession; all which they refer to one Name, and consider as one Thing. Hence it follows that when I examine by my other Senses a Thing I have seen, it is not in order to understand better the same Object which I had perceived by Sight, the Object of one Sense not being perceived by the other Senses. And when I look through a Microscope, it is not that I may perceive more clearly what I perceived already with my bare Eyes, the Object perceived by the Glass being quite different from the former. But in both cases my Aim is only to know what Ideas are connected together; and the more a Man knows of the Connexion of Ideas, the more he is said to know of the Nature of Things. What therefore if our Ideas are variable; what if our Senses are not in all Circumstances affected with the same Appearances? It will not thence follow, they are not to be trusted, or that they are inconsistent either with themselves or any thing else, except it be with your preconceived Notion of (I know not what) one single, unchanged, inperceivable, real Nature, marked by each Name: Which Prejudice seems to have taken its Rise from not rightly understanding the common Language of Men speaking of several distinct Ideas, as united into one thing by the Mind. And indeed there is Cause to suspect several erroneous Conceits of the Philosophers are owing to the same Original: While they began to build their Schemes, not so much on Notions as Words, which were framed by the Vulgar, merely for Conveniency and Dispatch in the common Actions of Life, without any regard to Speculation.

Hyl. Methinks I apprehend your Meaning.

Phil. It is your Opinion, the Ideas we perceive by our Senses are not real Things, but Images, or Copies of them. Our Knowledge therefore is no farther real, than as our Ideas are the true Representations of those Originals. But as these supposed Originals are in themselves unknown, it is impossible to know how far our Ideas resemble them; or whether they resemble them at all. We cannot therefore be sure we have any real Knowledge. Farther, as our Ideas are perpetually varied, without any Change in the supposed real Things, it necessarily follows they cannot all be true Copies of them: Or if some are, and others are not, it is impossible to distinguish the former from the latter. And this plunges us yet deeper in Uncertainty. Again, when we consider the Point, we cannot conceive how any Idea, or any thing like an Idea, should have an absolute Existence out of a Mind: Nor consequently, according to you, how there should be any real thing in Nature. The Result of all which is, that we are thrown into the most hopeless and abandoned Scepticism. Now give me leave to ask you, First, Whether your referring Ideas to certain absolutely existing unperceived Substances, as their Originals, be not the source of all this Scepticism? Secondly, Whether you are informed, either by Sense or Reason, of the Existence of those unknown Originals? And in case you are not, Whether it be not absurd to suppose them? Thirdly, Whether, upon Inquiry, you find there is any thing distinctly conceived or meant by the absolute or external Existence of unperceiving Substances? Lastly, Whether the premises considered, it be not the wisest way to follow Nature, trust your Senses, and laying aside all anxious Thought about unknown Natures or Substances, admit with the Vulgar those for real Things, which are perceived by the Senses?

Hyl. For the present, I have no Inclination to the answering Part. I would much rather see how you can get over what follows. Pray are not the Objects perceived by the Senses of one, likewise perceivable to others present? If there were an hundred more here, they would all see the Garden, the Trees, and Flowers as I see them. But they are not in the same manner affected with the Ideas I frame in my Imagination. Does not this make a Difference between the former sort of Objects and the latter?

Phil. I grant it does. Nor have I ever denied a Difference between the Objects of Sense and those of Imagination. But what would you infer from thence? You cannot say that sensible Objects exist unperceived, because they are perceived by many.

Hyl. I own, I can make nothing of that Objection: But it hath led me into another. Is it not your opinion that by our Senses we perceive only the Ideas existing in our Minds?

Phil. It is.

Hyl. But the same Idea which is in my Mind, cannot be in yours, or in any other Mind. Doth it not therefore follow from your Principles, that no Two can see the same thing? And is not this highly absurd?

Phil. If the Term same be taken in the vulgar Acceptation, it is certain, (and not at all repugnant to the Principles I maintain) that different Persons may perceive the same Thing; or the same Thing or Idea exist in different Minds. Words are of arbitrary Imposition; and since Men are used to apply the Word same where no Distinction or Variety is perceived, and I do not pretend to alter their Perceptions, it follows, that as Men have said before, several saw the same thing, so they may upon like Occasions still continue to use the same Phrase, without any Deviation either from Propriety of Language, or the Truth of Things. But if the Term same be used in the Acceptation of Philosophers, who pretend to an abstracted Notion of Identity, then, according to their sundry Definitions of this Notion, (for it is not yet agreed wherein that Philosophic Identity consists) it may or may not be possible for divers Persons to perceive the same thing. But whether Philosophers shall think fit to call a thing the same or no, is, I conceive, of small Importance. Let us suppose several Men together, all endued with the same Faculties, and consequently affected in like sort by their Senses, and who had yet never known the Use of Language; they would without question agree in their Perceptions. Though perhaps, when they came to the Use of Speech, some regarding the Uniformness of what was perceived, might call it the same thing: Others especially regarding the Diversity of Persons who perceived, might choose the Denomination of different things. But who sees not that all the Dispute is about a Word? to wit, Whether what is perceived by different Persons, may yet have the Term same applied to it? Or suppose a House, whose Walls or outward Shell remaining unaltered, the Chambers are all pulled down, and new ones built in their place; and that you should call this the same, and I should say it was not the same House: Would we not for all this perfectly agree in our Thoughts of the House, considered in it self? And would not all the Difference consist in a Sound? If you should say, We differ in our Notions; for that you superadded to your Idea of the House the simple abstracted Idea of Identity, whereas I did not; I would tell you I know not what you mean by that abstracted Idea of Identity; and should desire you to look into your own Thoughts, and be sure you understood your self.——Why so silent, Hylas? Are you not yet satisfied, Men may dispute about Identity and Diversity, without any real Difference in their Thoughts and Opinions, abstracted from Names? Take this farther Reflexion with you: That whether Matter be allowed to exist or no, the Case is exactly the same as to the Point in hand. For the Materialists themselves acknowledge what we immediately perceive by our Senses, to be our own Ideas. Your Difficulty therefore, that no two see the same thing, makes equally against the Materialists and me.

Hyl. But they suppose an external Archetype, to which referring their several Ideas, they may truly be said to perceive the same thing.

Phil. And (not to mention your having discarded those Archetypes) so may you suppose an external Archetype on my Principles; external, I mean, to your own Mind; though indeed it must be supposed to exist in that Mind which comprehends all things; but then this serves all the Ends of Identity, as well as if it existed out of a Mind. And I am sure you yourself will not say It is less intelligible.

Hyl. You have indeed clearly satisfied me, either that there is no Difficulty at bottom in this Point: or if there be, that it makes equally against both Opinions.

Phil. But that which makes equally against two contradictory Opinions, can be a Proof against neither.

Hyl. I acknowledge it. But after all, Philonous, when I consider the Substance of what you advance against Scepticism, it amounts to no more than this. We are sure that we really see, hear, feel; in a word, that we are affected with sensible Impressions.

Phil. And how are we concerned any farther? I see this Cherry. I feel it, I taste it: and I am sure nothing cannot be seen, or felt, or tasted: It is therefore real. Take away the Sensations of Softness, Moisture, Redness, Tartness, and you take away the Cherry. Since it is not a Being distinct from Sensations; a Cherry, I say, is nothing but a Congeries of sensible Impressions, or Ideas perceived by various Senses: Which Ideas are united into one thing (or have one Name given them) by the Mind; because they are observed to attend each other. Thus when the Palate is affected with such a particular Taste, the Sight is affected with a red Colour, the Touch with Roundness, Softness, &c. Hence, when I see, and feel, and taste, in sundry certain manners, I am sure the Cherry exists, or is real; its Reality being in my Opinion nothing abstracted from those Sensations. But if by the Word Cherry you mean an unknown Nature distinct from all those sensible Qualities, and by its Existence something distinct from its being perceived; then indeed I own, neither you nor I, nor any one else can be sure it exists.

Hyl. But what would you say, Philonous, if I should bring the very same Reasons against the Existence of sensible Things in a Mind, which you have offered against their existing in a material Substratum?

Phil. When I see your Reasons, you shall hear what I have to say to them.

Hyl. Is the Mind extended or unextended?

Phil. Unextended, without doubt.

Hyl. Do you say the Things you perceive are in your Mind?

Phil. They are.

Hyl. Again, have I not heard you speak of sensible Impressions?

Phil. I believe you may.

Hyl. Explain to me now, O Philonous! how it is possible there should be room for all those Trees and Houses to exist in your Mind. Can extended Things be contained in that which is unextended? Or are we to imagine Impressions made on a Thing void of all Solidity? You cannot say Objects are in your Mind, as Books in your Study: Or that Things are imprinted on it, as the Figure of a Seal upon Wax. In what Sense therefore are we to understand those Expressions? Explain me this if you can: And I shall then be able to answer all those Queries you formerly put to me about my Substratum.

Phil. Look you, Hylas, when I speak of Objects as existing in the Mind or imprinted on the Senses; I would not be understood in the gross literal Sense, as when Bodies are said to exist in a place, or a Seal to make an Impression upon Wax. My Meaning is only that the Mind comprehends or perceives them; and that it is affected from without, or by some Being distinct from itself. This is my Explication of your Difficulty; and how it can serve to make your Tenet of an unperceiving material Substratum intelligible, I would fain know.

Hyl. Nay, if that be all, I confess I do not see what Use can be made of it. But are you not guilty of some Abuse of Language in this?

Phil. None at all: It is no more than common Custom, which you know is the Rule of Language, hath authorized: Nothing being more usual, than for Philosophers to speak of the immediate Objects of the Understanding as Things existing in the Mind. Nor is there any thing in this, but what is conformable to the general Analogy of Language; most part of the mental Operations being signified by Words borrowed from sensible Things; as is plain in the Terms Comprehend, Reflect, Discourse, &c. which being applied to the Mind, must not be taken in their gross original Sense.

Hyl. You have, I own, satisfied me in this Point: But there still remains one great Difficulty, which I know not how you will get over. And indeed it is of such Importance, that if you could solve all others, without being able to find a Solution for this, you must never expect to make me a Proselyte to your Principles.

Phil. Let me know this mighty Difficulty.

Hyl. The Scripture Account of the Creation, is what appears to me utterly irreconcileable with your Notions. Moses tells us of a Creation: A Creation of what? of Ideas? No certainly, but of Things, of real Things, solid corporeal Substances. Bring your Principles to agree with this, and I shall perhaps agree with you.

Phil. Moses mentions the Sun, Moon, and Stars, Earth and Sea, Plants and Animals: That all these do really exist, and were in the Beginning created by God, I make no question. If by Ideas, you mean Fictions and Fancies of the Mind, then these are no Ideas. If by Ideas, you mean immediate Objects of the Understanding, or sensible Things which cannot exist unperceived, or out of a Mind, then these Things are Ideas. But whether you do, or do not call them Ideas, it matters little. The Difference is only about a Name. And whether that Name be retained or rejected, the Sense, the Truth and Reality of Things continues the same. In common Talk, the Objects of our Senses are not termed Ideas but Things. Call them so still, Provided you do not attribute to them any absolute external Existence, and I shall never quarrel with you for a Word. The Creation therefore I allow to have been a Creation of Things, of Real Things. Neither is this in the least inconsistent with my Principles, as is evident from what I have now said; and would have been evident to you without this, if you had not forgotten what had been so often said before. But as for solid corporeal Substances, I desire you to shew where Moses makes any mention of them; and if they should be mentioned by him, or any other inspired Writer, it would still be incumbent on you to shew those Words were not taken in the vulgar Acceptation, for things falling under our Senses, but in the Philosophic Acceptation, for Matter, or an unknown Quiddity, with an absolute Existence. When you have proved these Points, then (and not till then) may you bring the Authority of Moses into our Dispute.

Hyl. It is in vain to dispute about a Point so clear. I am content to refer it to your own Conscience. Are you not satisfied there is some peculiar Repugnancy between the Mosaic Account of the Creation, and your Notions?

Phil. If all possible Sense, which can be put on the first Chapter of Genesis, may be conceived as consistently with my Principles as any other, then it has no peculiar Repugnancy with them. But there is no Sense you may not as well conceive, believing as I do. Since, beside Spirits, all you conceive are Ideas; and the Existence of these I do not deny. Neither do you pretend they exist without the Mind.

Hyl. Pray let me see any Sense you can understand it in.

Phil. Why, I imagine that if I had been present at the Creation, I should have seen Things produced into Being; that is, become perceptible, in the Order prescribed by the Sacred Historian. I ever before believed the Mosaic Account of the Creation, and now find no Alteration in my Manner of believing it. When Things are said to being or end their Existence, we do not mean this with regard to God, but His Creatures. All Objects are eternally known by God, or which is the same thing, have an Eternal Existence in his Mind: But when Things before imperceptible to Creatures, are by a Decree of God made perceptible to them; then are they said to begin a relative Existence, with respect to created Minds. Upon reading therefore the Mosaic Account of the Creation, I understand that the several Parts of the World became gradually perceiveable to finite Spirits, endowed with proper Faculties; so that whoever such were present, they were in truth perceived by them. This is the literal obvious Sense suggested to me, by the Words of the Holy Scripture: In which is included no Mention or no Thought, either of Substratum, Instrument, Occasion, or absolute Existence. And upon Inquiry, I doubt not, it will be found, that most plain honest Men, who believe the Creation, never think of those things any more than I. What metaphysical Sense you may understand it in, you only can tell.

Hyl. But, Philonous, you do not seem to be aware, that you allow created Things in the Beginning, only a relative, and consequently hypothetical Being: That is to say, upon Supposition there were Men to perceive them, without which they have no Actuality of absolute Existence, wherein Creation might terminate. Is it not therefore according to you plainly impossible, the Creation of any inanimate Creatures should precede that of Man? And is not this directly contrary to the Mosaic Account?

Phil. In Answer to that I say, First, Created Beings might begin to exist in the Mind of other created Intelligences, beside Men. You will not therefore be able to prove any Contradiction between Moses and my Notions, unless you first shew, there was no other Order of finite created Spirits in Being before Man. I say farther, in case we conceive the Creation, as we should at this time a Parcel of Plants and Vegetables of all sorts, produced by an invisible Power, in a Desert where no body was present: That this Way of explaining it or conceiving it, is consistent with my Principles, since they deprive you of nothing, either sensible or imaginable: That it exactly suits with the common, natural, and undebauched Notions of Mankind: That it manifests the Dependence of all Things on God; and consequently hath all the good Effect or Influence, which it is possible that important Article of our Faith should have in making Men humble, thankful, and resigned to their Creator. I say moreover, that in this naked Conception of Things, divested of Words, there will not be found any Notion of what you call the Actuality of absolute Existence. You may indeed raise a Dust with those Terms, and so lengthen our Dispute to no purpose. But I intreat you calmly to look into your own Thoughts, and then tell me if they are not a useless and unintelligible Jargon.

Hyl. I own, I have no very clear Notion annexed to them. But what say you to this? Do you not make the Existence of sensible Things consist in their being in a Mind? And were not all Things eternally in the Mind of God? Did they not therefore exist from all Eternity, according to you? And how could that which was Eternal, be created in Time? Can any thing be clearer or better connected than this?

Phil. And are you not too of Opinion, that God knew all Things from Eternity?

Hyl. I am.

Phil. Consequently they always had a Being in the Divine Intellect.

Hyl. This I acknowledge.

Phil. By your own Confession therefore, nothing is New, or begins to be, in respect of the Mind of God. So we are agreed in that Point.

Hyl. What shall we make then of the Creation?

Phil. May we not understand it to have been intirely in respect of finite Spirits; so that Things, with regard to us, may properly be said to begin their Existence, or be created, when God decreed they should become perceptible to intelligent Creatures, in that Order and Manner which He then established, and we now call the Laws of Nature? You may call this a relative, or hypothetical Existence if you please. But so long as it supplies us with the most natural, obvious, and literal Sense of the Mosaic History of the Creation; so long as it answers all the religious Ends of that great Article; in a word, so long as you can assign no other Sense or Meaning in its stead; why should we reject this? Is it to comply with a ridiculous Sceptical Humour of making every thing Nonsense and Unintelligible? I am sure you cannot say, it is for the Glory of God. For allowing it to be a thing possible and conceivable, that the corporeal World should have an absolute Subsistence extrinsical to the Mind of God, as well as to the Minds of all created Spirits: Yet how could this set forth either the Immensity or Omniscience of the Deity, or the necessary and immediate Dependence of all things on Him? Nay, would it not rather seem to derogate from those Attributes?

Hyl. Well, but as to this Decree of God’s, for making Things perceptible: What say you, Philonous, is it not plain, God did either execute that Decree from all Eternity, or at some certain time began to will what He had not actually willed before, but only designed to will. If the former, then there could be no Creation or Beginning of Existence in finite Things. If the latter, then we must acknowledge something new to befal the Deity; which implies a sort of Change: and all Change argues Imperfection.

Phil. Pray consider what you are doing. Is it not evident, this Objection concludes equally against a Creation in any Sense; nay, against every other Act of the Deity, discoverable by the Light of Nature? None of which can we conceive, otherwise than as performed in Time, and having a Beginning. God is a Being of transcendent and unlimited Perfections: His Nature therefore is incomprehensible to finite Spirits. It is not therefore to be expected, that any Man, whether Materialist or Immaterialist, should have exactly just Notions of the Deity, His Attributes, and ways of Operation. If then you would infer any thing against me, your Difficulty must not be drawn from the Inadequateness of our Conceptions of the Divine Nature, which is unavoidable on any Scheme; but from the Denial of Matter, of which there is not one Word, directly or indirectly, in what you have now objected.

Hyl. I must acknowledge, the Difficulties you are concerned to clear, are such only as arise from the Non-existence of Matter, and are peculiar to that Notion. So far you are in the right. But I cannot by any means bring my self to think there is no such peculiar Repugnancy between the Creation and your Opinion: though indeed where to fix it, I do not distinctly know.

Phil. What would you have! do I not acknowledge a twofold State of Things, the one Ectypal or Natural, the other Archetypal and Eternal? The former was created in Time; the latter existed from Everlasting in the Mind of God. Is not this agreeable to the common Notions of Divines? or is any more than this necessary in order to conceive the Creation? But you suspect some peculiar Repugnancy, though you know not where it lies. To take away all Possibility of Scruple in the case, do but consider this one Point. Either you are not able to conceive the Creation on any Hypothesis whatsoever; and if so, there is no ground for Dislike or Complaint against any particular Opinion on that Score: Or you are able to conceive it; and if so, why not on my Principles, since thereby nothing conceivable is taken away? You have all along been allowed the full Scope of Sense, Imagination, and Reason. Whatever therefore you could before apprehend, either immediately or mediately by your Senses, or by Ratiocination from your Senses; whatever you could perceive, imagine or understand, remains still with you. If therefore the Notion you have of the Creation by other Principles be intelligible, you have it still upon mine; if it be not intelligible, I conceive it to be no Notion at all; and so there is no Loss of it. And indeed it seems to me very plain, that the Supposition of Matter, that is, a thing perfectly unknown and inconceivable, cannot serve to make us conceive any thing. And I hope, it need not be proved to you, that if the Existence of Matter doth not make the Creation conceivable, the Creation’s being without it inconceivable, can be no Objection against its Non-Existence.

Hyl. I confess, Philonous, you have almost satisfied me in this Point of the Creation. Phil. I would fain know why you are not quite satisfied. You tell me indeed of a

Repugnancy between the Mosaic History and Immaterialism: But you know not where it lies. Is this reasonable, Hylas? Can you expect I should solve a Difficulty without knowing what it is? But to pass by all that, would not a Man think you were assured there is no Repugnancy between the received Notions of Materialists and the inspired Writings?

Hyl. And so I am.

Phil. Ought the Historical Part of Scripture to be understood in a plain obvious Sense, or in a Sense which is metaphysical, and out of the way?

Hyl. In the plain Sense, doubtless.

Phil. When Moses speaks of Herbs, Earth, Water, &c. as having been created by God; think you not the sensible Things, commonly signified by those Words, are suggested to every unphilosophical Reader?

Hyl. I cannot help thinking so.

Phil. And are not all Ideas, or Things perceived by Sense, to be denied a real Existence by the Doctrine of the Materialists?

Hyl. This I have already acknowledged.

Phil. The Creation therefore, according to them, was not the Creation of Things sensible, which have only a relative Being, but of certain unknown Natures, which have an absolute Being, wherein Creation might terminate.

Hyl. True.

Phil. Is it not therefore evident, the Asserters of Matter destroy the plain obvious Sense of Moses, with which their Notions are utterly inconsistent; and instead of it obtrude on us I know not what, something equally unintelligible to themselves and me?

Hyl. I cannot contradict you.

Phil. Moses tells us of a Creation. A Creation of what? of unknown Quiddities, of Occasions, or Substratums? No certainly; but of Things obvious to the Senses. You must first reconcile this with your Notions, if you expect I should be reconciled to them.

Hyl. I see you can assault me with my own Weapons.

Phil. Then as to absolute Existence, was there ever known a more jejeune Notion than that? Something it is, so abstracted and unintelligible, that you have frankly owned you could not conceive it, much less explain any thing by it. But allowing Matter to exist, and the Notion of absolute Existence to be as clear as Light; yet was this ever known to make the Creation more credible? Nay hath it not furnished the Atheists and Infidels of all Ages, with the most plausible Argument against a Creation? That a corporeal Substance, which hath an absolute Existence without the Minds of Spirits, should be produced out of nothing by the mere Will of a Spirit, hath been looked upon as a thing so contrary to all Reason, so impossible and absurd, that not only the most celebrated among the Ancients, but even divers Modern and Christian Philosophers have thought Matter coeternal with the Deity. Lay these things together, and then judge you whether Materialism disposes Men to believe the Creation of Things.

Hyl. I own, Philonous, I think it does not. This of the Creation is the last Objection I can think of; and I must needs own it hath been sufficiently answered as well as the rest. Nothing now remains to be overcome, but a sort of unaccountable Backwardness that I find in my self towards your Notions.

Phil. When a Man is swayed, he knows not why, to one Side of the Question; Can this, think you, be any thing else but the Effect of Prejudice, which never fails to attend old and rooted Notions? And indeed in this respect I cannot deny the Belief of Matter to have very much the Advantage over the contrary Opinion, with Men of a learned Education.

Hyl. I confess it seems to be as you say.

Phil. As a Balance therefore to this Weight of Prejudice, let us throw into the Scale the great Advantages that arise from the Belief of Immaterialism, both in regard to Religion and Humane Learning. The Being of a God, and Incorruptibility of the Soul, those great Articles of Religion, are they not proved with the clearest and most immediate Evidence? When I say the Being of a God, I do not mean an obscure general Cause of Things, whereof we have no Conception, but God, in the strict and proper Sense of the Word. A Being whose Spirituality, Omnipresence, Providence, Omniscience, Infinite Power and Goodness, are as conspicuous as the Existence of sensible Things, of which (notwithstanding the fallacious Pretences and affected Scruples of Scepticks) there is no more reason to doubt, than of our own Being. Then with relation to Humane Sciences; in Natural Philosophy, what Intricacies, what Obscurities, what Contradictions, hath the Belief of Matter led Men into! To say nothing of the numberless Disputes about its Extent, Continuity, Homogeneity, Gravity, Divisibility, &c. do they not pretend to explain all things by Bodies operating on Bodies, according to the Laws of Motion? and yet, are they able to comprehend how one Body should move another? Nay, admitting there was no Difficulty in reconciling the Notion of an inert Being with a Cause; or in conceiving how an Accident might pass from one Body to another; yet by all their strained Thoughts and extravagant Suppositions, have they been able to reach the mechanical Production of any one Animal or Vegetable Body? Can they account by the Laws of Motion, for Sounds, Tastes, Smells, or Colours, or for the regular Course of Things? Have they accounted by Physical Principles for the Aptitude and Contrivance, even of the most inconsiderable Parts of the Universe? But laying aside Matter and corporeal Causes, and admitting only the Efficiency of an All-perfect Mind, are not all the Effects of Nature easy and intelligible? If the Phænomema are nothing else but Ideas; God is a Spirit, but Matter an unintelligent, unperceiving Being. If they demonstrate an unlimited Power in their Cause; God is Active and Omnipotent, but Matter an inert Mass. If the Order, Regularity, and Usefulness of them, can never be sufficiently admired; God is infinitely Wise and Provident, but Matter destitute of all Contrivance and Design. These surely are great Advantages in Physics. Not to mention that the Apprehension of a distant Deity, naturally disposes Men to a Negligence in their moral Actions, which they would be more cautious of, in case they thought Him immediately present, and acting on their Minds without the Interposition of Matter, or unthinking Second Causes. Then in Metaphysics; what Difficulties concerning Entity in Abstract, Substantial Forms, Hylarchic Principles, Plastic Natures, Substance and Accident, Principle of Individuation, Possibility of Matter’s thinking, Origin of Ideas, the Manner how two independent Substances, so widely different as Spirit and Matter, should mutually operate on each other? What Difficulties, I say, and endless Disquisitions concerning these and innumerable other the like Points, do we escape by supposing only Spirits and Ideas? Even the Mathematicks themselves, if we take away the absolute Existence of extended Things, become much more clear and easy; the most shocking Paradoxes and intricate Speculations in those Sciences, depending on the infinite Divisibility of finite Extension, which depends on that Supposition. But what need is there to insist on the particular Sciences? Is not that Opposition to all Science whatsoever, that Phrensy of the ancient and modern Scepticks, built on the same Foundation? Or can you produce so much as one Argument against the Reality of corporeal Things, or in behalf of that avowed utter Ignorance of their Natures, which doth not suppose their Reality to consist in an external absolute Existence? Upon this Supposition indeed, the Objections from the Change of Colours in a Pigeon’s Neck, or the Appearance of the broken Oar in the Water, must be allowed to have Weight. But those and the like Objections vanish, if we do not maintain the Being of absolute external Originals, but place the Reality of Things in Ideas, fleeting indeed, and changeable; however not changed at random, but according to the fixed Order of Nature. For herein consists that Constancy and Truth of Things, which secures all the Concerns of Life, and distinguishes that which is real from the irregular Visions of the Fancy.

Hyl. I agree to all you have now said, and must own that nothing can incline me to embrace your Opinion, more than the Advantages I see it is attended with. I am by Nature lazy; and this would be a mighty Abridgment in Knowledge. What Doubts, what Hypotheses, what Labyrinths of Amusement, what Fields of Disputation, what an Ocean of false Learning, may be avoided by that single Notion of Immaterialism?

Phil. After all, is there any thing farther remaining to be done? You may remember you promised to embrace that Opinion, which upon Examination should appear most agreeable to common Sense, and remote from Scepticism. This by your own Confession is that which denies Matter, or the absolute Existence of corporeal Things. Nor is this all; The same Notion has been proved several Ways, viewed in different Lights, pursued in its Consequences, and all Objections against it cleared. Can there be a greater Evidence of its Truth? or is it possible it should have all the Marks of a true Opinion and yet be false?

Hyl. I own my self intirely satisfied for the present in all respects. But what Security can I have that I shall still continue the same full Assent to your Opinion, and that no unthought-of Objection or Difficulty will occur hereafter?

Phil. Pray, Hylas, do you in other Cases, when a Point is once evidently proved, withhold your Assent on account of Objections or Difficulties it may be liable to? Are the Difficulties that attend the Doctrine of incommensurable Quantities, of the Angle of Contact, of the Asymptotes to Curves or the like, sufficient to make you hold out against Mathematical Demonstration? Or will you disbelieve the Providence of God, because there may be some particular things which you know not how to reconcile with it? If there are Difficulties attending Immaterialism, there are at the same time direct and evident Proofs of it. But for the Existence of Matter, there is not one Proof, and far more numerous and insurmountable Objections lie against it. But where are those might Difficulties you insist on? Alas! you know not where or what they are; something which may possibly occur hereafter. If this be a sufficient Pretence for withholding your full Assent, you should never yield it to any Proposition, how free soever from Exceptions, how clearly and solidly soever demonstrated.

Hyl. You have satisfied me, Philonous.

Phil. But to arm you against all future Objections, do but consider, That which bears equally hard on two contradictory Opinions, can be Proof against neither. Whenever therefore any Difficulty occurs, try if you can find a Solution for it on the Hypothesis of the Materialists. Be not deceived by Words; but sound your own Thoughts. And in case you cannot conceive it easier by the help of Materialism, it is plain it can be no Objection against Immaterialism. Had you proceeded all along by this Rule, you would probably have spared yourself abundance of trouble in objecting; since of your Difficulties I challenge you to shew one that is explained by Matter: nay, which is not more unintelligible with, than without that Supposition, and consequently makes rather against than for it. You should consider in each Particular, whether the Difficulty arises from the Non-existence of Matter. If it doth not, you might as well argue from the infinite Divisibility of Extension against the Divine Prescience, as from such a Difficulty against Immaterialism. And yet upon Recollection I believe you will find this to have been often, if not always the Case. You should likewise take heed not to argue on a petitio Principii. One is apt to say, The unknown Substances ought to be esteemed real Things, rather than the Ideas in our Minds: And who can tell but the unthinking external Substance may concur as a Cause or Instrument in the Production of our Ideas? But is not this proceeding on a Supposition that there are such external Substances? And to suppose this, is it not begging the Question? But above all things you should beware of imposing on your self by that vulgar Sophism, which is called Ignoratio Elenchi. You talked often as if you thought I maintained the Non-existence of sensible Things: Whereas in truth no one can be more thorowly assured of their Existence than I am: And it is you who doubt; I should have said, positively deny it. Every thing that is seen, felt, heard, or any way perceived by the Senses, is on the Principles I embrace, a real Being, but not on yours. Remember, the Matter you contend for is an unknown somewhat, (if indeed it may be termed somewhat) which is quite stripped of all sensible Qualities, and can neither be perceived by Sense, nor apprehended by the Mind. Remember, I say, that it is not any Object which is hard or soft, hot or cold, blue or white, round or square, &c. For all these things I affirm do exist. Though indeed I deny they have an Existence distinct from being perceived; or that they exist out of all Minds whatsoever. Think on these Points; let them be attentively considered and still kept in view. Otherwise you will not comprehend the state of the Question; without which your Objections will always be wide of the Mark, and instead of mine, may possibly be directed (as more than once they have been) against your own Notions.

Hyl. I must needs own, Philonous, nothing seems to have kept me from agreeing with you more than this same mistaking the Question. In denying Matter, at first glimpse I am tempted to imagine you deny the things we see and feel: but upon Reflexion find there is no Ground for it. What think you therefore of retaining the Name Matter, and applying it to sensible Things? This may be done without any Change in your Sentiments: And believe me it would be a Means of reconciling them to some Persons, who may be more shocked at an Innovation in Words than in Opinion.

Phil. With all my heart: Retain the Word Matter, and apply it to the Objects of Sense, if you please, provided you do not attribute to them any Subsistence distinct from their being perceived. I shall never quarrel with you for an Expression. Matter, or material Substance, are Terms introduced by Philosophers; and as used by them, imply a sort of Independency, or a Subsistence distinct from being perceived by a Mind: But are never used by common People; or if ever, it is to signify the immediate Objects of Sense. One would think therefore, so long as the Names of all particular Things, with the Terms sensible, Substance, Body, Stuff, and the like, are retained, the Word Matter should be never missed in common Talk. And in Philosophical Discourses it seems the best way to leave it quite out; since there is not perhaps any one thing that hath more favoured and strengthned the depraved Bent of the Mind towards Atheism, than the Use of that general confused Term.

Hyl. Well but, Philonous, since I am content to give up the Notion of an unthinking Substance exterior to the Mind, I think you ought not to deny me the Privilege of using the Word Matter as I please, and annexing it to a Collection of sensible Qualities subsisting only in the Mind. I freely own there is no other Substance in a strict Sense, than Spirit. But I have been so long accustomed to the Term Matter, that I know not how to part with it. To say, There is no Matter in the World, is still shocking to me. Whereas to say, There is no Matter, if by that Term be meant an unthinking Substance existing without the Mind: But if by Matter is meant some sensible Thing, whose Existence consists in being perceived, then there is Matter: This Distinction gives it quite another Turn: And Men will come into your Notions with small Difficulty, when they are proposed in that manner. For after all, the Controversy about Matter in the strict Acceptation of it, lies altogether between you and the Philosophers; whose Principles, I acknowledge, are not near so natural, or so agreeable to the common Sense of Mankind, and Holy Scripture, as yours. There is nothing we either desire or shun, but as it makes, or is apprehended to make some Part of our Happiness or Misery. But what hath Happiness or Misery, Joy or Grief, Pleasure or Pain, to do with absolute Existence, or with unknown Entities, abstracted from all Relation to us? It is evident, Things regard us only as they are pleasing or displeasing: And they can please or displease, only so far forth as they are perceived. Farther therefore we are not concerned; and thus far you leave things as you found them. Yet still there is something new in this Doctrine. It is plain, I do not now think with the Philosophers, nor yet altogether with the Vulgar. I would know how the Case stands in that respect: Precisely, what you have added to, or altered in my former Notions.

Phil. I do not pretend to be a Setter-up of New Notions. My Endeavours tend only to unite and place in a clearer Light that Truth, which was before shared between the Vulgar and the Philosophers: the former being of Opinion, that those Things they immediately perceive are the real Things; and the latter, that the Things immediately perceived, are Ideas which exist only in the Mind. Which Two Notions put together, do in effect constitute the Substance of what I advance.

Hyl. I have been a long time distrusting my Senses; methought I saw things by a dim Light, and through false Glasses. Now the Glasses are removed, and a new Light breaks in upon my Understanding. I am clearly convinced that I see things in their native Forms; and am no longer in Pain about their unknown Natures or absolute Existence. This is the State I find my self in at present: Though indeed the Course that brought me to it, I do not yet thorowly comprehend. You set out upon the same Principles that Academicks, Cartesians, and the like Sects, usually do; and for a long time it looked as if you were advancing their Philosophical Scepticism; but in the End your Conclusions are directly opposite to theirs.

Phil. You see, Hylas, the Water of yonder Fountain, how it is forced upwards, in a round Column, to a certain Height; at which it breaks and falls back into the Bason from whence it rose: Its Ascent as well as Descent, proceeding from the same uniform Law or Principle of Gravitation. Just so, the same Principles which at first View lead to Scepticism, pursued to a certain Point, bring Men back to common Sense.

FINIS.